### v1

# Learning Domain-Invariant Features for Out-of-Context News Detection

Abstract.
Multimodal out-of-context news is a common type of misinformation on online media platforms. This involves posting a caption, alongside an invalid out-of-context news image. Reflecting its importance, researchers have developed models to detect such misinformation. However, a common limitation of these models is that they only consider the scenario where pre-labeled data is available for each domain, failing to address the out-of-context news detection on unlabeled domains (e.g., unverified news on new topics or agencies). In this work, we therefore focus on domain adaptive out-of-context news detection. In order to effectively adapt the detection model to unlabeled news topics or agencies, we propose ConDA-TTA (Contrastive Domain Adaptation with Test-Time Adaptation) which applies contrastive learning and maximum mean discrepancy (MMD) to learn the domain-invariant feature. In addition, it leverages target domain statistics during test-time to further assist domain adaptation. Experimental results show that our approach outperforms baselines in 5 out of 7 domain adaptation settings on two public datasets, by as much as 2.93% in F1 and 2.08% in accuracy.
contrastive learning, domain adaptation, out-of-context news detection, multimodal learning
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ACM Conference XXXX; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/24/10ccs: Computing methodologies Natural language processingccs: Information systems Web applicationsccs: Computing methodologies Computer vision
1. Introduction
Online news platforms suffer from the release and spread of misinformation. Disseminating misinformation is especially concerning in news topics related to key societal issues, such as public health and politics, where it could result in misleading public awareness and arousing public panic. A common (Fazio, 2020) yet subtle type of misinformation is image repurposing, also called out-of-context news. Unlike DeepFake (Dolhansky et al., 2020) models which generate non-existing images, this approach attaches a false claim with a real image outside of its original context (see Figure 1). This lowers the technical threshold (Fazio, 2020) of producing misinformation since anyone can trivially attach a fabricated caption to an image and post it online, thereby enlarging its scope of occurrence. Complicating matters, out-of-context news cannot be detected using a sole modality input (as in isolation, each modality may be valid).
Figure 1. Examples of out-of-context news of three different news topics from the Twitter-COMMs dataset.
Out-of-context news detection has received a growing attention in recent years. Existing works (Biamby et al., 2022; Luo et al., 2021; Abdelnabi et al., 2022; Mu et al., 2023) fine-tune pre-trained vision and text encoders to represent the image and text, and then devise neural networks to measure vision-text consistency. More recent works (Dai et al., 2024; Zhang et al., 2023) query the Multimodal Large Language Model (MLLM) to obtain the prediction with explanations. Nevertheless, their approaches do not address the adaptation to entirely new topics or news agencies. Due to the wide array of news topics shared online, it is infeasible to re-train the detector every time a new topic emerges. Therefore, a detection model that can adapt to new topics will greatly relieve the operational burden. Additionally, considering annotation costs, it is equally critical to build a detection model that is able to adapt to other news agencies with minimal efforts, so that the model developed for one news agency can be easily extended to other popular news agencies, e.g. USA Today. Because of this, we argue that it is important to devise domain adaptive out-of-context news detection model.
Unfortunately, prior works neglect addressing the domain adaption issue on out-of-context news detection. Trained on sufficient labeled data from each domain, existing models tend to capture domain-specific knowledge which is not shared among different domains. Although this knowledge is helpful for improving the model’s performance on annotated domains, it results in inferior performance on unannotated domains (Wang et al., 2018). In light of this, we believe that learning domain-invariant features will help the detection model adapt to unlabeled domains. Though important, enabling domain adaptive capabilities on this task comes with several challenges. First, existing out-of-context news datasets only cover a small number of topics/agencies. This makes it challenging for the detection model to learn domain-invariant features, because the learned features might be biased towards the limited number of source domains. Second, in (social media) news dataset, the data distributions of different domains vary, since news from different domains tend to have different writing styles and different image styles. This makes domain-invariant knowledge learning challenging. It is critical to ensure that transferable domain-invariant knowledge is actually learned. Otherwise, the model may bypass learning transferable domain knowledge and end up learning semantic patterns as a shortcut to fulfill the classification task (Li et al., 2023a; Chi et al., 2023).
To tackle the aforementioned challenges, we propose Contrastive Domain Adaptation with Test-Time Adaptation (ConDA-TTA) for domain adaptive out-of-context news detection. ConDA-TTA first uses an MLLM to directly encode the image and text into a multimodal feature representation. After feature encoding, we adopt the contrastive loss to learn a representation in the projected space where the original news stays further away from the out-of-context news. To overcome the shortage of labeled source domains, we use unlabeled target domain data as well, to learn less biased domain-invariant features. We further apply Maximum Mean Discrepancy (MMD) to measure and reduce the discrepancy of the learned representations between the source and target domain. In this way, MMD ensures that the domain-invariant features are captured from the learned representations. Finally, we incorporate Test-Time Adaptation (TTA) to update the statistic-related model parameters in the evaluation phase so that the model can better adapt to the target domain.
To summarize, our main contributions are as follows:
  • To the best of our knowledge, we are the first to investigate domain adaptation in out-of-context news detection. To address the challenges, we propose a novel approach using contrastive domain adaptation with test-time adaptation (ConDA-TTA).
  • We evaluate our approach and demonstrate that it achieves state-of-the-art (SOTA) performance under most domain adaptation settings. Notably, it outperforms the baseline by as much as 2.93% in F1 and 2.08% in accuracy when the domain is defined as news topic. Additionally, it outperforms the baseline by as much as 1.82% in F1 and 1.84% in accuracy when the domain is defined as news agency.
  • We conduct a comprehensive ablation study and show that MMD tends to be the most contributing component to the domain adaptation on the Twitter-COMMs dataset, while TTA tends to be the most contributing component on the NewsCLIPpings dataset.
2. Related Work
2.1. Out-of-Context News Detection
Prior works on out-of-context news detection mainly adopt three approaches: (i) fine-tuning or (ii) prompting large vision and language models, and (iii) leveraging synthetic information. Biamby et al. (2022) fine-tunes CLIP (Radford et al., 2021) and a linear classifier to detect out-of-context news. Similarly, Luo et al. (2021) fine-tunes both CLIP and VisualBert (Li et al., 2019) with a linear classifier. Abdelnabi et al. (2022) concatenates the retrieved textual evidence with visual evidence and then inputs the learned representations into a simple classifier. These works involve fine-tuning a large multimodal pre-trained model (e.g. CLIP), which is very compute-intensive (e.g.  (Biamby et al., 2022) uses 8 days of training on 10 GPUs), and does not enable domain adaptation.
Later work leverages synthetic multimodal information. Shalabi et al. (2023) uses a stable diffusion model (Dhariwal and Nichol, 2021) to generate augmented image from text, and a pre-trained image captioning model to generate augmented text from image. (Yuan et al., 2023) extracts stances from the external multimodal evidences and uses it to enhance evidence-based out-of-context detection. More recent works focus on enhancing the interpretability. (Qi et al., 2024) adopts two-stage instruction tuning on Instruct-BLIP (Dai et al., 2024) to provide accurate and persuasive explanations to the prediction. (Zhang et al., 2023) uses neural symbolic model to enhance model interpretability. (Shalabi et al., 2024) fine-tunes the MiniGPT-4 to detect out-of-context news. Although effective, the above works do not address the domain adaptation problem in out-of-context news detection.
2.2. Domain Adaptive Fake News Detection
To the best of our knowledge, no prior work has explored domain adaptive out-of-context news detection. However, some prior works have investigated techniques to enable domain adaptation in multimodal fake news detection — a task close to out-of-context news detection. Some works (Wang et al., 2018; Zhang et al., 2020, 2021; Yuan et al., 2021) adopt adversarial learning through a domain classifier/discriminator to learn the domain-invariant feature. On text-based domain adaptation in fake news detection, Nan et al. (2022) adopts meta learning and Ding et al. (2022) further combines it with adversarial learning. Furthermore, incorporating user comments and user-news interaction information, Mosallanezhad et al. (2022) adopts reinforcement learning that exploits cross-domain and within-domain knowledge to achieve robustness in the target domain. Yue et al. (2022) proposes a contrastive domain adaptation method in order to reduce the intra-class discrepancy and enlarge the inter-class discrepancy. Lin et al. (2023) employs prompt engineering to learn language-agnostic contextual representations and models the domain-invariant structural features from the propagation threads. In contrast, our work focuses on the domain adaptation in out-of-context news detection and adopts both contrastive learning and test-time adaptation to address the unique challenges (Section  1) of this task.
2.3. Test-Time Adaptation
Test-time adaptation (TTA) is an emerging learning paradigm that increases the trained model’s generalization ability to the test samples under distribution shifts (Liang et al., 2023). We later employ this to improve our performance. Many works focus on adapting batch normalization (BN) layer’s parameters during model testing: Li et al. (2016) updates the BN statistics; Wang et al. (2020) further updates the learnable parameters using information entropy; Zhao et al. (2023) proposes to use exponential moving average (EMA) to update BN parameters and to address the imbalanced classes problem. Wang et al. (2022) proposes continual test-time adaptation for reducing error accumulation and catastrophic forgetting. Recent works propose to conduct sample filtering during TTA: Niu et al. (2022) filters out less reliable samples and maintains sample diversity at the same time; Niu et al. (2023) removes partial noisy samples with large gradient and encourages model weights to go to a flat minimum to achieve stable TTA. In our work, we combine test-time adaptation with contrastive learning, in order to further facilitate domain adaptation in out-of-context news detection.
3. Problem Statement
In out-of-context news detection, given the news image-text pair $(x_{img},x_{txt})$, the detection model is expected to predict whether it is falsified, i.e. out-of-context. If the image and text are not from the same news post, this image-text pair is considered as falsified (labeled as True), otherwise not falsified (labeled as False).
In this work, we tackle the challenge of domain adaptive out-of-context news detection. During training, we have access to the labeled source domain data and unlabeled target domain data. During testing, we evaluate our model’s performance on the labeled target domain data. The source domain and target domain are mutually exclusive. The target domain training data and testing data are mutually exclusive as well. Concretely, the dataset is comprised of news from $N$ domains. Each time we select $M$ domains as the target domain, $\mathcal{D}^{T}=\{\mathcal{D}^{T_{m}}\}_{m=1}^{M}$, and the rest $N-M$ domains as the source domain, $\mathcal{D}^{S}=\{\mathcal{D}^{S_{n}}\}_{n=1}^{N-M}$. The $n$-th source domain $\mathcal{D}^{S_{n}}$ contains labeled data, $\mathcal{D}^{S_{n}}=\{(x_{i}^{S_{n}},y_{i}^{S_{n}})\}_{i=1}^{|S_{n}|}$, where $x^{S_{n}}_{i}=(x_{img},x_{txt})_{i}^{S_{n}}$ denotes the $i$-th news image-text pair within domain $\mathcal{D}^{S_{n}}$, and $y_{i}^{S_{n}}$ denotes whether this image-text pair is falsified or not (True or False). In contrast, the $m$-th target domain contains unlabeled data: $\mathcal{D}^{T_{m}}=\{x_{j}^{T_{m}}\}_{j=1}^{|T_{m}|}$, where $x_{j}^{T_{m}}=(x_{img},x_{txt})_{j}^{T_{m}}$.
For simplicity, in the following notations, we use $T$ to denote target domain and $S$ to denote source domain.
4. Our Approach: ConDA-TTA
In this section, we describe our proposed approach for domain adaptive out-of-context news detection: Contrastive Domain Adaptation with Test-Time Adaptation (ConDA-TTA). The overall architecture of ConDA-TTA is illustrated in Figure 2. It consists of three components: (i) Multimodal Feature Encoder, which encodes the image and text pair into a multimodal representation; (ii) Domain-Invariant Feature Learning, which applies contrastive learning and maximum mean discrepancy (MMD) to learn the domain-invariant feature; and (iii) Test-Time Adaptation, which leverages unlabeled test set statistics to further adapt the model to the target domain. Finally, we discuss the training details of our approach.
Figure 2. The overall model architecture of ConDA-TTA. We first use the (i) Multimodal Feature Encoder to encode out-of-context news and its augmentation into multimodal representations. Then in the (ii) Domain-Invariant Feature Learning, we apply contrastive learning and maximum mean discrepancy (MMD) to learn the domain-invariant feature shared by different domains. Finally, we adopt the (iii) Test-Time Adaptation to update statistic-related model parameters in the evaluation phase.
4.1. Multimodal Feature Encoder
In order to capture the semantic nuance between original news and out-of-context news, we use MLLM to directly encode the news text and image into a meaningful multimodal representation.
Formally, the multimodal feature encoder is defined as follows:
(1) $\textbf{x}^{S}=\mathrm{MLLM}((x_{img},x_{txt})^{S}),\quad\textbf{x}^{T}=\mathrm{MLLM}((x_{img},x_{txt})^{T}),$
where $\textbf{x}^{S}$ and $\textbf{x}^{T}$ respectively denote the multimodal feature representation of news image-text pair $(x_{img},x_{txt})$ from the source and target domain.
4.2. Domain-Invariant Feature Learning
After encoding the news, we introduce a domain-invariant feature learning module to learn the domain-invariant feature from the news’ multimodal representation. Specifically, this module adopts contrastive learning to learn a more separable representation space and adopts MMD to capture the invariant features among different domains.
We first use a projection head to project the multimodal feature representation into a lower dimensional space. The projection head is composed of two linear layers, and can be expressed as follows:
(2) $\textbf{z}^{S}=\mathrm{Projection}(\textbf{x}^{S}),\quad\textbf{z}^{T}=\mathrm{Projection}(\textbf{x}^{T}).$
Next, we aim to train the projection head so that it can capture the domain-invariant features from $\textbf{x}^{S}$ and $\textbf{x}^{T}$. This is enabled by applying contrastive learning and the maximum mean discrepancy. Note, no label information is required within this module.
Contrastive Learning: Inspired by (Bhattacharjee et al., 2023), we leverage contrastive learning to learn more separable representations of the input multimodal features in the projected space. In this way, we expect that the learned representations of similar semantic meanings stay closer, and the learned representations of dissimilar semantic meanings stay further apart.
To achieve this, we adopt the contrastive loss used in  (Bhattacharjee et al., 2023). Applying contrastive loss pulls closer the positive pairs and pushes further apart the negative pairs. We augment each news item in order to construct the positive and negative pairs later. For each news item $\textbf{x}_{i}^{S}$ in a source domain training batch $\{\textbf{x}_{i}^{S}\}_{i=1}^{b}$, we generate one augmentation of it and encode it by the same multimodal feature encoder to obtain its representation, denoted as $\textbf{x}_{i+}^{S}$. In this way, we obtain $2|b|$ news items in the augmented batch. After projecting the input multimodal feature to the lower dimensional space, the positive pair is formed by the data item and its corresponding augmentation item, denoted as $(\textbf{z}_{i}^{S},\textbf{z}_{i+}^{S})$; the negative pairs are formed by the data item and the rest of the $2(|b|-1)$ items within the augmented batch, denoted as $(\textbf{z}_{i}^{S},\textbf{z}_{k}^{S})$. The contrastive loss is expressed as follows:
(3) $\mathcal{L}_{ctr}^{S}=-\sum_{(i,i_{+})\in b}\log\frac{\exp(\mathrm{sim}(\textbf{z}_{i}^{S},\textbf{z}_{i_{+}}^{S})/t)}{\sum_{k=1,k\neq i}^{2|b|}\exp(\mathrm{sim}(\textbf{z}_{i}^{S},\textbf{z}_{k}^{S})/t)}.$
Here, $b$ denotes the current batch and $|b|$ denotes the size of the current batch. $S$ denotes the source domain. $\textbf{z}_{i}^{S}$ and $\textbf{z}_{j}^{S}$ denote the learned representations of the $i$-th data item and its augmentation in the projected space. $t$ denotes the temperature coefficient. $\mathrm{sim}(\cdot)$ denotes the similarity metric. Here, cosine similarity is being used. The data augmentation on the target domain training batch and the contrastive loss computation for the target domain $\mathcal{L}_{ctr}^{T}$ is applied in the same way.
The performance of contrastive learning is largely influenced by the transformation used to generate positive samples (Tian et al., 2020). Following the data augmentation experiments in (Bhattacharjee et al., 2023), we include text-level: {synonym replacement, random swap, random crop}, and image-level: {random resize and crop, random horizontal flip, Gaussian blur} into the set of candidate data augmentation techniques (detailed descriptions can be found in the supplementary material). We also include the combination of text-level and image-level augmentations as the candidates. Our experiments show that all the above strategies result in similar performances (in terms of accuracy on the target domain), but applying Gaussian blur to the image results in the best overall performance (averaged over all domain adaptation settings) on the target domain. Thus, in our approach, we apply Gaussian blur to the image while keeping the text unaltered to generate the positive sample of the anchor item. Note, the augmentation preserves the label.
Maximum Mean Discrepancy (MMD): Now that we have learned more separable representations, $\textbf{z}^{S}$ and $\textbf{z}^{T}$, in the projected space, we want to further ensure that these representations capture domain-invariant feature. To achieve this, we adopt maximum mean discrepancy (MMD) (Gretton et al., 2012). MMD is a kernel-based statistical test used to determine if two samples are drawn from different distributions (Gretton et al., 2012). We use it as a loss function to measure the discrepancy between two distributions. Here, we regard the learned representations $\textbf{z}^{S}$ and $\textbf{z}^{T}$ as the samples from two random variables $Z^{S}$ and $Z^{T}$, and we want their probability distributions to have as small as possible discrepancy. In this way, the distribution of $Z^{S}$ and $Z^{T}$ conditions less on a specific domain. Alternatively speaking, if MMD is small enough, we regard that it removes the domain-specific features from z and keeps the domain-invariant features.
The squared MMD between $Z^{S}$ and $Z^{T}$ in the reproducing kernel Hilbert space (RKHS) $\mathcal{H}$ is formulated as follows:
(4) $\mathrm{MMD}^{2}[\mathcal{F},Z^{S},Z^{T}]$ $=\left[\sup_{||f||_{\mathcal{H}}\leq 1}(\mathbf{E}[f(Z^{S})]-\mathbf{E}[f(Z^{T})])\right]^{2}$
$=||\mu_{S}-\mu_{T}||^{2}_{\mathcal{H}},$
where $\mu_{S}$ and $\mu_{T}$ are mean embeddings of $f(Z^{S})$ and $f(Z^{T})$, $f\in\mathcal{F}$. $\mathcal{F}$ is the unit ball function class in the Hilbert space $\mathcal{H}$.
Following prior work (Pan et al., 2010; Long et al., 2015; Bhattacharjee et al., 2023), we compute the MMD between features in the lower dimensional space (i.e. the projected z space). In our implementation, we use the empirical MMD (Gretton et al., 2012), which is expressed as follows:
(5) $\mathrm{MMD}[\mathcal{F},Z^{S},Z^{T}]=$ $\left[\frac{1}{m^{2}}\sum_{i,j=1}^{m}k(\textbf{z}_{i}^{S},\textbf{z}_{j}^{S})-\frac{2}{mn}\sum_{i,j=1}^{m,n}k(\textbf{z}_{i}^{S},\textbf{z}_{j}^{T})\right.$
$\left.+\frac{1}{n^{2}}\sum_{i,j=1}^{n}k(\textbf{z}_{i}^{T},\textbf{z}_{j}^{T})\right]^{\frac{1}{2}}.$
Here, both $m$ and $n$ are equal to the batch size. $k(\cdot,\cdot)$ denotes the kernel function. $\sigma$ is a free parameter. We use the Radial basis function kernel shown in the following:
(6) $k(\textbf{z}_{i},\textbf{z}_{j})=\exp\left(-\frac{||\textbf{z}_{i}-\textbf{z}_{j}||^{2}_{2}}{2\sigma^{2}}\right).$
Training: The learned domain-invariant representation, z, is then input to the classifier to get the predicted label $\hat{y}$ for the corresponding out-of-context news:
(7) $\hat{y}=\mathrm{CLS}(\textbf{z}).$
The loss function for the classifier is the cross-entropy loss, where $\hat{y}$ is the predicted label and $y$ is the ground true label:
(8) $\mathcal{L}_{CE}=-\mathbf{E}_{y\sim Y}\left[y\log(\hat{y})+(1-y)\log(1-\hat{y})\right].$
During training, both the projection head and the classifier have learnable parameters. The total loss is composed of (i) the contrastive losses from both the source and target data in the training set, (ii) the MMD between the source and target projected representations, and (iii) the cross-entropy losses of the predicted label for both the source data ($\mathcal{L}_{CE}$) and its augmentation ($\mathcal{L}_{CE+}$). This can be expressed as follows:
(9) $\mathcal{L}=\frac{1}{2}\lambda_{CE}(\mathcal{L}_{CE}^{S}+\mathcal{L}_{CE+}^{S})+\frac{1}{2}\lambda_{ctr}(\mathcal{L}_{ctr}^{S}+\mathcal{L}_{ctr}^{T})+\lambda_{MMD}\mathrm{MMD},$
where $\lambda_{CE}$ is the weight of the cross entropy losses, $\lambda_{ctr}$ is the weight of the contrastive losses and $\lambda_{MMD}$ is the weight of MMD.
4.3. Test-Time Adaptation
Domain-invariant features help the model to generalize to the target domain. In this section, we introduce test-time adaptation which leverages target domain statistics to further adapt the model to the target domain during evaluation.
Prior work (Li et al., 2016) observes that the statistics of the Batch Normalization (BN) layer contain domain-specific information. Concretely, they find that running the test set through the trained model achieves deep adaptation effects on the target domain. The reason is that this process allows BN statistics to keep track of the target domain statistics, which is conducive to the domain adaptation. As such, we introduce BN layers into the classifier.
At each training step, BN statistics compute and update the running estimates $\{\hat{\mu}_{k+1},\hat{\sigma}^{2}_{k+1}\}$ (Eq. (10)) based on its observed mean $\mu_{k}$ and variance $\sigma_{k}^{2}$ (Eq. (11)) of the layer batch inputs $\{x_{i}\}_{i=1}^{b}$:
(10) $\hat{\mu}_{k+1}=(1-\rho)\hat{\mu}_{k}+\rho\mu_{k},\quad\hat{\sigma}^{2}_{k+1}=(1-\rho)\hat{\sigma}^{2}_{k}+\rho\sigma^{2}_{k}$
(11) $\mu_{k}=\frac{1}{b}\sum_{i}x_{i},\quad\sigma^{2}_{k}=\frac{1}{b}\sum_{i}(x_{i}-\mu_{k})^{2}$
Here, $k$ denotes the $k$-th training step, and $\rho$ is the momentum with a default value of 0.1. BN statistics are then fixed (denoted as $\mu$ and $\sigma^{2}$) and used for normalization when the model is set to evaluation mode (pyt, 2024). During evaluation, we apply test-time adaptation to calibrate the BN statistics, i.e. the mean ($\mu$) and variance ($\sigma^{2}$). The BN layer is formulated as follows:
(12) $x_{out}=\frac{x_{in}-\mu}{\sqrt{\sigma^{2}+\epsilon}}*\gamma+\beta$
where $x_{in}$ is the BN layer input, $x_{out}$ is the output. $\gamma$ and $\beta$ are BN learnable parameters. $\epsilon$ is added for numerical stability.
During model training, since only the source domain data is passed through the classifier, BN statistics are computed based on the source domain statistics. During model evaluation, the target domain data is passed through the classifier. To integrate target domain statistics into the BN statistics, after each training epoch, we pass the unlabeled test set through the classifier. After that, we set the classifier to evaluation mode and evaluate it on the test set. In this way, the running estimates would keep the target domain statistics as well into record, which is helpful for the domain adaptation.
5. Experimental Design
5.1. Datasets
We conduct experiments on Twitter-COMMs and NewsCLIPpings, which are the only two out-of-context news datasets that conform to our problem definition (Section 3).
Twitter-COMMs (Biamby et al., 2022): This dataset is collected from Twitter solely on three topics: Covid-19 (Cv), Climate Change (Cl) and Military Vehicles (M). Each original tweet is represented by an image and its associated caption (text). Given a caption, the authors create out-of-context news by either randomly choosing an image (random negative) or by choosing the image whose original associated text is semantically the closest (hard negative). After running the script provided by the authors to download tweets and removing tweets that has an unavailable image, we obtain 2,143,934 items (Cv: 1,387,043, Cl: 512,490, M: 244,401) in the training set, and 22,082 items (Cv: 6,456, Cl: 8,488, M: 7,138) in the test set. The labels have a balanced distribution. Hence, we address domain adaptation among different news topics in Twitter-COMMs.
NewsCLIPpings (Luo et al., 2021): This dataset is derived from the VisualNews dataset (Liu et al., 2020), which is a benchmark news image captioning dataset. The news is collected from four agencies, namely Guardian (G), BBC (B), Washington Post (W) and USA Today (U). The NewsCLIPpings dataset is originally partitioned into four splits in which ways the out-of-context news is constructed. However, the analysis on the underlying VisualNews dataset (Liu et al., 2020) reveals that there are inter correlations among data from different news agencies (e.g. , the image captioning model trained on the USA Today data achieves 3.7 in CIDEr score (Vedantam et al., 2015) when tested on the USA Today test set, yet only achieves 0.6 on the Guardian test set. ). Since NewsCLIPping is derived from VisualNews, we re-partition the dataset according to the news agencies. After downloading the dataset, we obtain a total number of 1,182,900 items (B: 205,968, G: 567,012, U: 228,393, W: 181,527) in the training set, and 124,168 items (B: 22,074, G: 59,163, U: 24,243, W: 18,688) in the test set. The labels have a balanced distribution. Hence, we address domain adaptation among different news agencies in NewsCLIPpings.
5.2. Implementation Details
Table 1. Evaluation of domain adaptation performance on TwitterCOMMs and NewsCLIPpings. Cv, Cl and M stand for covid-19, climate change and military vehicles respectively. B, G, U and W stand for BBC, Guardian, USA Today and Washington Post respectively. In X $\rightarrow$ Y, X denotes the source domain, Y denotes the target domain. The best performance is in bold text. The second best performance is underlined.
Model Twitter-COMMs NewsCLIPpings
Cv, Cl $\rightarrow$ M Cv, M $\rightarrow$ Cl Cl, M $\rightarrow$ Cv U, W $\rightarrow$ B U, W $\rightarrow$ G B, G $\rightarrow$ U B, G $\rightarrow$ W
F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc
Source only 70.00 72.36 78.48 78.83 76.92 77.87 67.19 67.90 75.08 75.23 79.77 79.94 78.60 78.67
Source and target 80.60 80.68 81.47 81.48 81.50 81.58 73.80 74.71 76.00 77.96 82.70 84.37 79.80 80.74
EANN 76.33 77.26 79.13 79.16 79.44 79.60 69.30 69.61 76.01 76.12 79.72 79.88 78.27 78.34
MDA-WS 76.83 77.09 76.12 76.99 78.39 78.55 69.62 69.79 74.88 74.89 79.54 79.73 78.22 78.23
CANMD 67.13 68.16 79.87 80.53 79.91 80.01 68.42 68.62 74.75 75.64 79.30 79.69 79.93 78.59
REAL-FND 73.50 74.82 79.13 79.36 76.64 77.62 69.70 69.96 75.20 75.30 79.95 80.00 78.50 78.52
ConDA-TTA 79.26 79.34 80.80 80.80 80.06 80.08 71.52 71.80 75.70 75.86 80.77 80.84 77.17 77.31
Multimodal Feature Encoder Selection: To find the optimal MLLM feature extractor, we experiment with (i) BLIP-2 (Li et al., 2023b) multimodal and (ii) unimodal feature extractor, (iii) CLIP (Radford et al., 2021) (only has unimodal feature extractors) and (iv) ALBEF (Li et al., 2021b) multimodal feature extractor. For (ii) and (iii), we experiment with vector concatenation and element-wise multiplication to obtain the multimodal feature vector.
We then train a one-layer linear classifier on the four features respectively and evaluate the performances. Experimental results show that BLIP-2 multimodal features gives the best performance (in terms of accuracy) on both the training and test sets. Note, we also evaluate the zero-shot performance of the BLIP-2 Image Text Matching head which only outputs a similarity score rather than the feature vector. We find it underperforms (around 0.5 in accuracy on both training and test set) using BLIP-2 multimodal feature directly.
Architecture: We use BLIP-2’s blip2_feature_extractor with the multimodal mode to embed the image and text, and obtain the multimodal embedding x of size 768. The projection head is composed of two linear layers with 768 and 500 (experimented with {300, 500, 768}) neurons respectively. The classifier is composed of three linear layers with 768, 768 and 2 neurons respectively, and two batch normalization layers following the first two linear layers. We use Tanh() as the activation function. We apply dropout both before the first linear layer and the third linear layer. The dropout rate is set to 0.2 (we experimented with {0.2, 0.5} and without dropout).
Training: We perform training using the Adam optimizer with a batch size of 256 for 20 epochs. The early stopping epoch is set to 5. The learning rate is set to 2e-4 for the Twitter-COMMs dataset and 1e-4 for the NewsCLIPpings dataset. For the loss function, $\lambda_{MMD}$ is set to 1 and both $\lambda_{CE}$ and $\lambda_{ctr}$ are set to 0.5.
[…]
5.3. Baseline Models
Considering that no prior work has looked into domain adaptive out-of-context news detection, we select SOTA baseline models from a similar task: domain adaptive fake news detection. For fair comparison, we use BLIP-2’s blip2_feature_extractor with the multimodal mode to encode the image and text into a multimodal feature vector.
Source only serves as the naive baseline for the chosen backbone feature encoder. It uses one linear layer and the softmax function as the classifier. During training, it only uses the source domain.
Source and target uses the same architecture as source only. The difference is that, during training, it uses all domains, including both the source and target domains. Therefore, it can be viewed as the upper bound performance.
EANN (Wang et al., 2018) is a multimodal fake news detection model that uses adversarial learning to derive event-invariant features which benefit fake news detection on newly arrived events. Since the original EANN implementation encodes the image and text modality separately, we modify it to accommodate multimodal inputs.
MDA-WS (Li et al., 2021a) is an early fake news detection framework based on multi-source domain adaptation and weak supervision. It uses adversarial learning to learn the domain-invariant feature, and incorporates prior knowledge to assign pseudo labels to target domain news for weak supervision.
CANMD (Yue et al., 2022) is a domain adaptive fake news detection model. It firstly pre-trains the model on the source training data and then uses the pseudo labels generated from it on the target training data to adapt the model. During adaptation, it aims to reduce intra-class discrepancy and enlarge inter-class discrepancy.
REAL-FND (Mosallanezhad et al., 2022) is a reinforcement-learning (RL) based domain adaptive fake news detection model. It first obtains a reasonable news representation and then trains the fake news and domain classifiers. Using the representation as the state and the classifiers’ losses as the reward, it then trains an RL agent that learns a domain-invariant news representation.
6. Result and Analysis
Based on the above experimental setup, we evaluate our proposed ConDA-TTA and investigate the contribution of each model component afterwards. We also include TSNE visualization and parameter sensitivity analysis.
6.1. Experimental Result
Table 2. Evaluation results on ablating $\mathcal{L}_{ctr}$, MMD and TTA of ConDA-TTA. Cv, Cl and M stand for covid-19, climate change and military vehicles respectively. B, G, U and W stand for BBC, Guardian, USA Today and Washington Post respectively. In X $\rightarrow$ Y, X denotes the source domain, Y denotes the target domain. The best performance is in bold text. The second best performance is underlined.
Model Twitter-COMMs NewsCLIPpings
Cv, Cl $\rightarrow$ M Cv, M $\rightarrow$ Cl Cl, M $\rightarrow$ Cv U, W $\rightarrow$ B U, W $\rightarrow$ G B, G $\rightarrow$ U B, G $\rightarrow$ W
F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc
w/o $\mathcal{L}_{ctr}$ 78.07 78.12 80.59 80.59 80.10 80.11 70.80 70.97 75.07 75.28 80.27 80.33 77.33 77.50
w/o MMD 75.85 75.92 80.47 80.48 79.14 79.17 71.38 71.46 75.64 75.80 79.93 79.98 77.22 77.35
w/o TTA 78.58 78.59 80.73 80.73 80.52 80.58 70.20 70.26 75.68 75.68 80.75 80.77 77.89 77.89
ConDA-TTA 79.26 79.34 80.80 80.80 80.06 80.08 71.52 71.80 75.70 75.86 80.77 80.84 77.17 77.31
Table 1 shows the evaluation performance of ConDA-TTA and the baseline models. We use F1 and accuracy (Acc) as the evaluation metrics.
Naive Baseline and Upper Bound: We first present our observations on the naive baseline (source only in Table 1) and the supervised training (source and target in Table 1) results. In our experiments, we find that using BLIP-2’s multimodal representation with a plain linear classifier already achieves strong performances (around 70 - 80% zero-shot accuracy under most domain adaptation scenarios) on this task. Nevertheless, we observe around 2-10% performance gap between the naive baseline and the supervised training result. Considering that BLIP-2 has 2.7b parameters, it is not trivial to apply fine-tuning to improve its domain adaptation performance. This suggests the need for developing domain adaptation techniques on top of it to mitigate the gap.
For the news topics in Twitter-COMMs, adapting to Military Vehicles (see Section 5.1) has the largest domain adaptation gap (10.6% in F1 and 8.32% in accuracy), whilst adapting to Climate Change and Covid-19 has a relatively smaller domain adaptation gap (2.99% in F1 and 2.65% in accuracy, 4.58% in F1 and 3.71% in accuracy). We conjecture that the differences are potentially caused by the different news styles. We manually check a batch of 10 randomly sampled news from each topic. We find that Climate Change and Covid-19 tend to include images with more abundant information, including things such as embedded texts, numbers, and even news screenshots. However, for the Military Vehicles domain, the news’ image usually contains one clear subject: the military vehicle itself. At the same time, it requires more specialized knowledge to determine whether the image pairs the text within this topic.
For the news agencies in NewsCLIPpings, adapting to BBC and USA Today has the largest domain adaptation gap (6.61% in F1 and 6.81% in accuracy, 2.93% in F1 and 4.43% in accuracy), while adapting to the Guardian and the Washington Post has a relatively smaller gap. Although the gap is not consistently significant, it can affect heavily if deployed in a news platforms where a large amount of news is posted everyday. Thus, we believe that leveraging these powerful MLLM feature encoders and further mitigating its domain adaptation gap with a small training cost are crucial to achieve promising performance on unlabeled, newly emergent domains.
Baseline Comparison: We next compare our model’s performance with the baseline models. On both datasets, our proposed ConDA-TTA outperforms the baseline models across most metrics. Notably, on the Twitter-COMMs dataset, ConDA-TTA outperforms the best baseline model by 2.93% in F1 and by 2.08% in accuracy when the source domains are Covid-19 (Cv) and Climate Change (Cl), and the target domain is Military Vehicles (M). Specifically, we observe negative transfer for CANMD under this setting. Moreover, ConDA-TTA achieves 80.80% on both F1 and accuracy when the source domains are Covid-19 and Military Vehicles, and the target domain is Climate Change, surpassing the baseline performances by 0.93% and 0.27% respectively. Additionally, it achieves 80.06% and 80.08% on F1 and accuracy when the source domains are Climate Change and Military Vehicles, and the target domain is Covid-19.
On the NewsCLIPpings dataset, when the source domains are USA Today (U) and Washington Post (W) and the target domain is BBC (B), our model outperforms the best baseline model by 1.84% in accuracy and 1.82% F1 accuracy. When the source domains are BBC (B) and Guardian (G) and the target domain is USA Today (U), our model outperforms the best baseline model by 0.82% in F1 and 0.84% in accuracy. When the source domains are USA Today and Washington Post, and the target domain is Guardian, our model achieves the second best performance in both metrics. However, we also observe that when the source domains are BBC and Guardian, and target domain is Washington Post, all models see slightly negative transfer in terms of accuracy. This is in-line with the observations (Section  5.1) on the underlying dataset (Liu et al., 2020), where the authors find that training their image captioning model on either the BBC or Guardian results in a comparatively low CIDEr score on the Washington Post. As such, we conjecture that the information contained in the BBC and Guardian is not enough for enhancing the model’s domain adaptation performance on the Washington Post.
[…]
6.2. Ablation Study
We next present an ablation study on the different components in ConDA-TTA. Specifically, we ablate $\mathcal{L}_{ctr}$, MMD and TTA from ConDA-TTA. The evaluation results are summarized in Table 2.
On the Twitter-COMMs dataset, we observe that ablating MMD causes the biggest performance drop in all three domain adaptation settings, respectively: 3.41% (F1) and 3.42% (Acc), 0.33% (F1) and 0.32% (Acc), 0.92% (F1) and 0.91% (Acc). This suggests that MMD is the most important component in ConDA-TTA. Additionally, ablating the contrastive loss results in the second biggest performance decrease when the target domain is Military Vehicles and Climate Change, respectively: 1.19% (F1) and 1.22% (Acc), 0.21% (F1) and 0.21% (Acc). However, we see a marginal performance increase (0.04% in F1 and 0.03% in Acc) when the target domain is Covid-19. Lastly, ablating TTA causes a 0.68% (F1) and 0.75% (Acc) performance drop when the target domain is Military Vehicles, and causes a 0.07% performance drop in both F1 and Acc when the target domain is Climate Change. When the target domain is Covid-19, we again see a slight increase (0.46% in F1 and 0.50% in Acc) in performance. Although ConDA-TTA (as a whole) improves the domain adaptation performance on Covid-19 (Section 6.1), not every component of it contributes under this setting.
On the NewsCLIPpings dataset, we observe similar degradation when ablating these three components when the target domain is the BBC, Guardian and USA Today. Specifically, when ablating TTA, the performance decreases (1.22% in F1 ad 1.53% in Acc) most significantly on BBC, while less significantly on Guardian (0.02% in F1 and 0.18% in Acc) and USA Today (0.02% in F1 and 0.07% in Acc). When ablating MMD, F1 drops by 0.04%, 0.06% and 0.84%, and the Acc drops by 0.04%, 0.06% and 0.86%. When ablating the contrastive loss, the F1 drops by 0.62%, 0.63% and 0.5%, and the Acc drops by 0.53%, 0.58% and 0.51%. However, when adapting to Washington Post, ablating the contrastive loss, MMD and TTA all bring slight performance increase (0.04% - 0.72%). We conjecture that this observation echos previous analysis (Section 6.1) that the information contained in BBC and Guardian is not enough for assisting the domain adaptation on Washington Post, making each component less effective.
In summary, the above results confirm the efficacy of the contrastive loss, MMD and TTA in most domain adaptation scenarios on both datasets. Concretely, MMD tends to contribute the most to the domain adaptation on Twitter-COMMs; TTA tends to contribute the most on NewsCLIPpings. However, we note that, in some cases, the techniques are less effective on the NewsCLIPpings dataset. We believe this is due to that under some settings, the source domain contains less abundant information for adaptation to the target domain (Section 6.1).
6.3. Visualization
To show that ConDA-TTA effectively learns the domain-invariant feature, we adopt TSNE (Van der Maaten and Hinton, 2008) to visualize the multimodal embeddings x (also the input embedding to ConDA-TTA) and the learned domain-invariant feature z projected to the 2-d space (Figure 3).
Figure 3. TSNE visualization of the multimodal feature x and the learned domain-invariant feature z under Cv, Cl $\rightarrow$ M.
Here we illustrate the TSNE visualization of setting Cv, Cl $\rightarrow$ M. We can tell from the figure that feature z from different topics blend better than x, where we see clearer boundaries between different topics. This suggests that our model has removed domain-specific features from x to some extent. This further shows that our approach is able to effectively learn domain-invariant feature, and transfers out-of-context news detection knowledge to the target domain (Table 1).
6.4. Sensitivity Analysis
The contrastive loss and the MMD are two crucial components in our approach. This section briefly studies how their weights, $\lambda_{ctr}$ and $\lambda_{MMD}$, in the loss function affect the performance of ConDA-TTA.
Figure 4. ConDA-TTA’s performance (in Acc) with different $\lambda_{MMD}$ and $\lambda_{ctr}$ values. The legend denotes the target domain.
Figure 4 illustrates ConDA-TTA’s performance with $\lambda_{MMD}=\{0.5,1,2,5\}$ and $\lambda_{ctr}=\{0.1,0.5,1,2\}$. When setting $\lambda_{MMD}$ to 1, our model achieves the best performance in 4 (M, Cl, B, U) out of 7 domain adaptation settings, and achieves close to best performance in the remaining 3 settings. Overall, $\lambda_{MMD}$ being too high causes the accuracy to drop a little. This might be because the model has put comparatively less emphasis on learning the detection of out-of-context news (i.e. the cross-entropy loss in Eq. (9) receives less attention). For $\lambda_{ctr}$, we find that setting it to 0.5 achieves the best performance in 5 settings. When the value goes up, the model performance tends to decrease gradually. We conjecture this is because the similarity (Section 4.2) among original news becomes smaller, making it slightly difficult to find the decision boundary.
7. Conclusion
This paper has proposed ConDA-TTA, a domain adaptive out-of-context news detection model that can adapt to both unlabeled news topics and news agencies. ConDA-TTA first encodes the news using the BLIP-2 multimodal feature encoder and then adopts contrastive learning and MMD to learn the domain-invariant features. During test-time, we further incorporate target domain statistics into the classifier of ConDA-TTA, in order to better adapt the model to the target domain. Experimental results show the efficacy of our approach. These advantages enable ConDA-TTA to transfer the learned out-of-context news detection knowledge to emerging events or news platforms, lessening the burden of online news content moderators.

### v2

@@
 Learning Domain-Invariant Features for Out-of-Context News Detection
-Abstract.
-Multimodal out-of-context news is a common type of misinformation on online media platforms.
-This involves posting a caption, alongside an invalid out-of-context news image.
-Reflecting its importance, researchers have developed models to detect such misinformation.
-However, a common limitation of these models is that they only consider the scenario where pre-labeled data is available for each domain, failing to address the out-of-context news detection on unlabeled domains (e.g., unverified news on new topics or agencies).
+Abstract
+Out-of-context news is a common type of misinformation on online media platforms.
+This involves posting a caption, alongside a mismatched news image.
+Existing out-of-context news detection models only consider the scenario where pre-labeled data is available for each domain, failing to address the out-of-context news detection on unlabeled domains (e.g.
+news topics or agencies).
 In this work, we therefore focus on domain adaptive out-of-context news detection.
-In order to effectively adapt the detection model to unlabeled news topics or agencies, we propose ConDA-TTA (Contrastive Domain Adaptation with Test-Time Adaptation) which applies contrastive learning and maximum mean discrepancy (MMD) to learn the domain-invariant feature.
-In addition, it leverages target domain statistics during test-time to further assist domain adaptation.
-Experimental results show that our approach outperforms baselines in 5 out of 7 domain adaptation settings on two public datasets, by as much as 2.93% in F1 and 2.08% in accuracy.
-contrastive learning, domain adaptation, out-of-context news detection, multimodal learning
-copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ACM Conference XXXX; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/24/10ccs: Computing methodologies Natural language processingccs: Information systems Web applicationsccs: Computing methodologies Computer vision
-1. Introduction
+In order to effectively adapt the detection model to unlabeled news topics or agencies, we propose ConDA-TTA (Contrastive Domain Adaptation with Test-Time Adaptation) which applies contrastive learning and maximum mean discrepancy (MMD) to learn domain-invariant features.
+In addition, we leverage test-time target domain statistics to further assist domain adaptation.
+Experimental results show that our approach outperforms baselines in most domain adaptation settings on two public datasets, by as much as 2.93% in F1 and 2.08% in accuracy.
+Learning Domain-Invariant Features for Out-of-Context News Detection
+Yimeng Gu1, Mengqi Zhang2, Ignacio Castro1, Shu Wu3, Gareth Tyson4 1Queen Mary University of London, 2Shandong University, 3CRIPC, Chinese Academy of Science, 4The Hong Kong University of Science and Technology (GZ) yimeng.gu@qmul.ac.uk
+1 Introduction
 Online news platforms suffer from the release and spread of misinformation.
-Disseminating misinformation is especially concerning in news topics related to key societal issues, such as public health and politics, where it could result in misleading public awareness and arousing public panic.
-A common (Fazio, 2020) yet subtle type of misinformation is image repurposing, also called out-of-context news.
+A common Fazio (2020) yet subtle type of misinformation is image repurposing, also called out-of-context news.
 Unlike DeepFake (Dolhansky et al., 2020) models which generate non-existing images, this approach attaches a false claim with a real image outside of its original context (see Figure 1).
-This lowers the technical threshold (Fazio, 2020) of producing misinformation since anyone can trivially attach a fabricated caption to an image and post it online, thereby enlarging its scope of occurrence.
-Complicating matters, out-of-context news cannot be detected using a sole modality input (as in isolation, each modality may be valid).
-Figure 1. Examples of out-of-context news of three different news topics from the Twitter-COMMs dataset.
+This lowers the technical threshold Fazio (2020) of producing misinformation since anyone can trivially attach a fabricated caption to an image and post it online, thereby enlarging its scope of occurrence.
+Figure 1: Examples of out-of-context news of three different news topics from the Twitter-COMMs dataset.
 Out-of-context news detection has received a growing attention in recent years.
-Existing works (Biamby et al., 2022; Luo et al., 2021; Abdelnabi et al., 2022; Mu et al., 2023) fine-tune pre-trained vision and text encoders to represent the image and text, and then devise neural networks to measure vision-text consistency.
-More recent works (Dai et al., 2024; Zhang et al., 2023) query the Multimodal Large Language Model (MLLM) to obtain the prediction with explanations.
-Nevertheless, their approaches do not address the adaptation to entirely new topics or news agencies.
+Existing works (Biamby et al., 2022; Luo et al., 2021; Abdelnabi et al., 2022; Mu et al., 2023) fine-tune pre-trained vision and text encoders to represent the image and text.
+More recent works Dai et al.
+(2024); Zhang et al.
+(2023) query Multimodal Large Language Models (MLLMs) to obtain the prediction with explanations.
+However, their approaches do not address the model adaptation to entirely new topics or news agencies.
 Due to the wide array of news topics shared online, it is infeasible to re-train the detector every time a new topic emerges.
-Therefore, a detection model that can adapt to new topics will greatly relieve the operational burden.
-Additionally, considering annotation costs, it is equally critical to build a detection model that is able to adapt to other news agencies with minimal efforts, so that the model developed for one news agency can be easily extended to other popular news agencies, e.g.
-USA Today.
-Because of this, we argue that it is important to devise domain adaptive out-of-context news detection model.
+Considering annotation costs, it is equally critical to build a detection model that can adapt to other news agencies with minimal efforts.
 Unfortunately, prior works neglect addressing the domain adaption issue on out-of-context news detection.
-Trained on sufficient labeled data from each domain, existing models tend to capture domain-specific knowledge which is not shared among different domains.
-Although this knowledge is helpful for improving the model’s performance on annotated domains, it results in inferior performance on unannotated domains (Wang et al., 2018).
+They tend to capture domain-specific knowledge not shared across different domains.
+Although this knowledge is helpful for improving the model’s performance on annotated domains, it results in inferior performance on unannotated domains Wang et al.
+(2018).
 In light of this, we believe that learning domain-invariant features will help the detection model adapt to unlabeled domains.
-Though important, enabling domain adaptive capabilities on this task comes with several challenges.
+However, this comes with several challenges.
 First, existing out-of-context news datasets only cover a small number of topics/agencies.
 This makes it challenging for the detection model to learn domain-invariant features, because the learned features might be biased towards the limited number of source domains.
-Second, in (social media) news dataset, the data distributions of different domains vary, since news from different domains tend to have different writing styles and different image styles.
+Second, in (social media) news dataset, the data distributions of different domains vary, since news from different domains tend to have different writing styles and image styles.
 This makes domain-invariant knowledge learning challenging.
 It is critical to ensure that transferable domain-invariant knowledge is actually learned.
-Otherwise, the model may bypass learning transferable domain knowledge and end up learning semantic patterns as a shortcut to fulfill the classification task (Li et al., 2023a; Chi et al., 2023).
+Otherwise, the model may bypass learning it and end up learning semantic patterns as a shortcut to fulfill the classification task (Li et al., 2023b; Chi et al., 2023).
 To tackle the aforementioned challenges, we propose Contrastive Domain Adaptation with Test-Time Adaptation (ConDA-TTA) for domain adaptive out-of-context news detection.
 ConDA-TTA first uses an MLLM to directly encode the image and text into a multimodal feature representation.
-After feature encoding, we adopt the contrastive loss to learn a representation in the projected space where the original news stays further away from the out-of-context news.
-To overcome the shortage of labeled source domains, we use unlabeled target domain data as well, to learn less biased domain-invariant features.
-We further apply Maximum Mean Discrepancy (MMD) to measure and reduce the discrepancy of the learned representations between the source and target domain.
-In this way, MMD ensures that the domain-invariant features are captured from the learned representations.
+After that, we adopt the contrastive loss to learn a representation in the projected space where the original news stays further away from the out-of-context news.
+To overcome the shortage of labeled source domains, we take advantage of the unlabeled target domain data as well, to learn less biased domain-invariant features.
+To ensure that domain-invariant features are captured from the learned representations, we then apply Maximum Mean Discrepancy (MMD) to reduce the discrepancy of the learned representations between the source and target domain.
 Finally, we incorporate Test-Time Adaptation (TTA) to update the statistic-related model parameters in the evaluation phase so that the model can better adapt to the target domain.
 To summarize, our main contributions are as follows:
-  • To the best of our knowledge, we are the first to investigate domain adaptation in out-of-context news detection.
-To address the challenges, we propose a novel approach using contrastive domain adaptation with test-time adaptation (ConDA-TTA).
-  • We evaluate our approach and demonstrate that it achieves state-of-the-art (SOTA) performance under most domain adaptation settings.
-Notably, it outperforms the baseline by as much as 2.93% in F1 and 2.08% in accuracy when the domain is defined as news topic.
-Additionally, it outperforms the baseline by as much as 1.82% in F1 and 1.84% in accuracy when the domain is defined as news agency.
-  • We conduct a comprehensive ablation study and show that MMD tends to be the most contributing component to the domain adaptation on the Twitter-COMMs dataset, while TTA tends to be the most contributing component on the NewsCLIPpings dataset.
-2. Related Work
-2.1. Out-of-Context News Detection
-Prior works on out-of-context news detection mainly adopt three approaches: (i) fine-tuning or (ii) prompting large vision and language models, and (iii) leveraging synthetic information.
+  • We are the first to investigate domain adaptation in out-of-context news detection.
+To address the challenges, we propose a novel approach that learns domain-invariant features through MMD and TTA.
+  • We evaluate our approach and demonstrate its effectiveness.
+Notably, it outperforms the baseline by as much as 2.93% in F1 when the domain is defined as news topic.
+Additionally, it outperforms the baseline by as much as 1.82% in F1 when the domain is defined as news agency.
+  • We conduct a comprehensive ablation study and show that MMD is the most contributing component to the domain adaptation on Twitter-COMMs, while TTA is the most contributing component on NewsCLIPpings.
+2 Related Work
+2.1 Out-of-Context News Detection
+Prior works on out-of-context news detection mainly adopt three technical routes: (i) fine-tuning or (ii) prompting large vision and language models, and (iii) leveraging synthetic information.
 Biamby et al.
-(2022) fine-tunes CLIP (Radford et al., 2021) and a linear classifier to detect out-of-context news.
+(2022) fine-tunes CLIP (Radford et al., 2021) to detect out-of-context news.
 Similarly, Luo et al.
-(2021) fine-tunes both CLIP and VisualBert (Li et al., 2019) with a linear classifier.
+(2021) fine-tunes both CLIP and VisualBert (Li et al., 2019).
 Abdelnabi et al.
-(2022) concatenates the retrieved textual evidence with visual evidence and then inputs the learned representations into a simple classifier.
-These works involve fine-tuning a large multimodal pre-trained model (e.g.
-CLIP), which is very compute-intensive (e.g.
-(Biamby et al., 2022) uses 8 days of training on 10 GPUs), and does not enable domain adaptation.
+(2022) additionally leverages the retrieved textual and visual evidences.
+These works involve fine-tuning large multimodal pre-trained models, which is very compute-intensive and does not enable domain adaptation.
 Later work leverages synthetic multimodal information.
 Shalabi et al.
-(2023) uses a stable diffusion model (Dhariwal and Nichol, 2021) to generate augmented image from text, and a pre-trained image captioning model to generate augmented text from image.
-(Yuan et al., 2023) extracts stances from the external multimodal evidences and uses it to enhance evidence-based out-of-context detection.
-More recent works focus on enhancing the interpretability.
-(Qi et al., 2024) adopts two-stage instruction tuning on Instruct-BLIP (Dai et al., 2024) to provide accurate and persuasive explanations to the prediction.
-(Zhang et al., 2023) uses neural symbolic model to enhance model interpretability.
-(Shalabi et al., 2024) fine-tunes the MiniGPT-4 to detect out-of-context news.
+(2023) generates augmented image from text, and generates augmented text from image.
+Yuan et al.
+(2023) extracts stances from external multimodal evidences to enhance the detection.
+More recent works focus on enhancing interpretability.
+Qi et al.
+(2024) adopts two-stage instruction tuning on Instruct-BLIP Dai et al.
+(2024) to provide accurate and persuasive explanations to the prediction.
+Zhang et al.
+(2023) uses neural symbolic model to enhance model interpretability.
+Shalabi et al.
+(2024) fine-tunes MiniGPT-4 to detect out-of-context news.
 Although effective, the above works do not address the domain adaptation problem in out-of-context news detection.
-2.2. Domain Adaptive Fake News Detection
+2.2 Domain Adaptive Fake News Detection
 To the best of our knowledge, no prior work has explored domain adaptive out-of-context news detection.
-However, some prior works have investigated techniques to enable domain adaptation in multimodal fake news detection — a task close to out-of-context news detection.
-Some works (Wang et al., 2018; Zhang et al., 2020, 2021; Yuan et al., 2021) adopt adversarial learning through a domain classifier/discriminator to learn the domain-invariant feature.
-On text-based domain adaptation in fake news detection, Nan et al.
-(2022) adopts meta learning and Ding et al.
-(2022) further combines it with adversarial learning.
-Furthermore, incorporating user comments and user-news interaction information, Mosallanezhad et al.
+However, existing works have investigated domain adaptation in fake news detection — a task close to out-of-context news detection.
+Many works Wang et al.
+(2018); Zhang et al.
+(2020, 2021); Yuan et al.
+(2021) adopt adversarial learning to learn the domain-invariant feature.
+Furthermore, by incorporating user comments and user-news interaction information, Mosallanezhad et al.
 (2022) adopts reinforcement learning that exploits cross-domain and within-domain knowledge to achieve robustness in the target domain.
@@
 In contrast, our work focuses on the domain adaptation in out-of-context news detection and adopts both contrastive learning and test-time adaptation to address the unique challenges (Section  1) of this task.
-2.3. Test-Time Adaptation
-Test-time adaptation (TTA) is an emerging learning paradigm that increases the trained model’s generalization ability to the test samples under distribution shifts (Liang et al., 2023).
-We later employ this to improve our performance.
-Many works focus on adapting batch normalization (BN) layer’s parameters during model testing: Li et al.
-(2016) updates the BN statistics; Wang et al.
-(2020) further updates the learnable parameters using information entropy; Zhao et al.
-(2023) proposes to use exponential moving average (EMA) to update BN parameters and to address the imbalanced classes problem.
-Wang et al.
-(2022) proposes continual test-time adaptation for reducing error accumulation and catastrophic forgetting.
-Recent works propose to conduct sample filtering during TTA: Niu et al.
-(2022) filters out less reliable samples and maintains sample diversity at the same time; Niu et al.
-(2023) removes partial noisy samples with large gradient and encourages model weights to go to a flat minimum to achieve stable TTA.
-In our work, we combine test-time adaptation with contrastive learning, in order to further facilitate domain adaptation in out-of-context news detection.
-3. Problem Statement
+3 Problem Statement
 In out-of-context news detection, given the news image-text pair $(x_{img},x_{txt})$, the detection model is expected to predict whether it is falsified, i.e.
 out-of-context.
-If the image and text are not from the same news post, this image-text pair is considered as falsified (labeled as True), otherwise not falsified (labeled as False).
-In this work, we tackle the challenge of domain adaptive out-of-context news detection.
+If the image and text are not from the same news post, this news is considered as falsified (labeled as True), otherwise not falsified (labeled as False).
+In this work, we tackle the challenges of domain adaptive out-of-context news detection.
 During training, we have access to the labeled source domain data and unlabeled target domain data.
@@
 The source domain and target domain are mutually exclusive.
-The target domain training data and testing data are mutually exclusive as well.
 Concretely, the dataset is comprised of news from $N$ domains.
 Each time we select $M$ domains as the target domain, $\mathcal{D}^{T}=\{\mathcal{D}^{T_{m}}\}_{m=1}^{M}$, and the rest $N-M$ domains as the source domain, $\mathcal{D}^{S}=\{\mathcal{D}^{S_{n}}\}_{n=1}^{N-M}$.
-The $n$-th source domain $\mathcal{D}^{S_{n}}$ contains labeled data, $\mathcal{D}^{S_{n}}=\{(x_{i}^{S_{n}},y_{i}^{S_{n}})\}_{i=1}^{|S_{n}|}$, where $x^{S_{n}}_{i}=(x_{img},x_{txt})_{i}^{S_{n}}$ denotes the $i$-th news image-text pair within domain $\mathcal{D}^{S_{n}}$, and $y_{i}^{S_{n}}$ denotes whether this image-text pair is falsified or not (True or False).
+The $n$-th source domain $\mathcal{D}^{S_{n}}$ contains labeled data, $\mathcal{D}^{S_{n}}=\{(x_{i}^{S_{n}},y_{i}^{S_{n}})\}_{i=1}^{|S_{n}|}$, where $x^{S_{n}}_{i}=(x_{img},x_{txt})_{i}^{S_{n}}$ denotes the $i$-th news image-text pair within domain $\mathcal{D}^{S_{n}}$, and $y_{i}^{S_{n}}$ denotes whether this image-text pair is falsified or not.
 In contrast, the $m$-th target domain contains unlabeled data: $\mathcal{D}^{T_{m}}=\{x_{j}^{T_{m}}\}_{j=1}^{|T_{m}|}$, where $x_{j}^{T_{m}}=(x_{img},x_{txt})_{j}^{T_{m}}$.
 For simplicity, in the following notations, we use $T$ to denote target domain and $S$ to denote source domain.
-4. Our Approach: ConDA-TTA
+4 Our Approach: ConDA-TTA
 In this section, we describe our proposed approach for domain adaptive out-of-context news detection: Contrastive Domain Adaptation with Test-Time Adaptation (ConDA-TTA).
 The overall architecture of ConDA-TTA is illustrated in Figure 2.
-It consists of three components: (i) Multimodal Feature Encoder, which encodes the image and text pair into a multimodal representation; (ii) Domain-Invariant Feature Learning, which applies contrastive learning and maximum mean discrepancy (MMD) to learn the domain-invariant feature; and (iii) Test-Time Adaptation, which leverages unlabeled test set statistics to further adapt the model to the target domain.
-Finally, we discuss the training details of our approach.
-Figure 2.
-The overall model architecture of ConDA-TTA.
-We first use the (i) Multimodal Feature Encoder to encode out-of-context news and its augmentation into multimodal representations.
-Then in the (ii) Domain-Invariant Feature Learning, we apply contrastive learning and maximum mean discrepancy (MMD) to learn the domain-invariant feature shared by different domains.
+It consists of three components: (i) Multimodal Feature Encoder, which encodes the image and text pair into a multimodal representation; (ii) Contrastive Domain Adaptation, which applies contrastive learning to learn a more separable representation space and maximum mean discrepancy (MMD) to learn the domain-invariant feature; and (iii) Test-Time Adaptation, which leverages unlabeled test set statistics to further adapt the model to the target domain.
+Figure 2: The model architecture of ConDA-TTA.
+We first use the (i) Multimodal Feature Encoder to encode the news and its augmentation into multimodal representations.
+Then in the (ii) Contrastive Domain Adpatation, we apply contrastive learning and maximum mean discrepancy (MMD) to learn the domain-invariant feature.
 Finally, we adopt the (iii) Test-Time Adaptation to update statistic-related model parameters in the evaluation phase.
-4.1. Multimodal Feature Encoder
-In order to capture the semantic nuance between original news and out-of-context news, we use MLLM to directly encode the news text and image into a meaningful multimodal representation.
+4.1 Multimodal Feature Encoder
+In order to capture the semantic differences between original news and out-of-context news, we use MLLM to directly encode the news text and image into a meaningful multimodal representation.
 Formally, the multimodal feature encoder is defined as follows:
-(1) $\textbf{x}^{S}=\mathrm{MLLM}((x_{img},x_{txt})^{S}),\quad\textbf{x}^{T}=\mathrm{MLLM}((x_{img},x_{txt})^{T}),$
-where $\textbf{x}^{S}$ and $\textbf{x}^{T}$ respectively denote the multimodal feature representation of news image-text pair $(x_{img},x_{txt})$ from the source and target domain.
-4.2. Domain-Invariant Feature Learning
-After encoding the news, we introduce a domain-invariant feature learning module to learn the domain-invariant feature from the news’ multimodal representation.
-Specifically, this module adopts contrastive learning to learn a more separable representation space and adopts MMD to capture the invariant features among different domains.
-We first use a projection head to project the multimodal feature representation into a lower dimensional space. The projection head is composed of two linear layers, and can be expressed as follows:
-(2) $\textbf{z}^{S}=\mathrm{Projection}(\textbf{x}^{S}),\quad\textbf{z}^{T}=\mathrm{Projection}(\textbf{x}^{T}).$
-Next, we aim to train the projection head so that it can capture the domain-invariant features from $\textbf{x}^{S}$ and $\textbf{x}^{T}$.
-This is enabled by applying contrastive learning and the maximum mean discrepancy.
+$\textbf{x}^{S}$ $=\mathrm{MLLM}((x_{img},x_{txt})^{S}),$ (1)
+$\textbf{x}^{T}$ $=\mathrm{MLLM}((x_{img},x_{txt})^{T}),$
+where $\textbf{x}^{S}$ and $\textbf{x}^{T}$ respectively denote the multimodal feature representation of the news image-text pair $(x_{img},x_{txt})$ from the source and target domain.
+4.2 Contrastive Domain Adaptation
+After encoding the news, we introduce the contrastive domain adaptation module.
+Specifically, this module first adopts contrastive learning to learn a more separable representation space for news image-text pairs and then uses MMD to capture the invariant features among different domains.
+We first use a projection head to project the multimodal feature representation into a lower dimensional space. The projection head can be expressed as follows:
+$\textbf{z}^{S}=\mathrm{Projection}(\textbf{x}^{S}),\quad\textbf{z}^{T}=\mathrm{Projection}(\textbf{x}^{T}).$ (2)
+Next, we aim to train the projection head so that it can learn a more separable representation space and capture the domain-invariant features from $\textbf{x}^{S}$ and $\textbf{x}^{T}$ afterwards.
+This is achieved by applying the contrastive learning and maximum mean discrepancy respectively.
 Note, no label information is required within this module.
-Contrastive Learning: Inspired by (Bhattacharjee et al., 2023), we leverage contrastive learning to learn more separable representations of the input multimodal features in the projected space.
-In this way, we expect that the learned representations of similar semantic meanings stay closer, and the learned representations of dissimilar semantic meanings stay further apart.
-To achieve this, we adopt the contrastive loss used in  (Bhattacharjee et al., 2023).
+Contrastive Learning: Contrastive learning has been shown effective in learning better representations for the classification task Qian et al.
+(2022).
+Inspired by Bhattacharjee et al.
+(2023), we leverage contrastive learning to learn more separable representations of the input multimodal features in the projected space.
+In this way, we expect that the learned representations of similar semantic meanings stay closer, and the learned representations of dissimilar semantic meanings stay further apart, so that it could make the classification easier.
+To achieve this, we adopt the contrastive loss used in  Bhattacharjee et al.
+(2023).
 Applying contrastive loss pulls closer the positive pairs and pushes further apart the negative pairs.
-We augment each news item in order to construct the positive and negative pairs later.
+To construct the positive and negative pairs, we first augment each news item.
 For each news item $\textbf{x}_{i}^{S}$ in a source domain training batch $\{\textbf{x}_{i}^{S}\}_{i=1}^{b}$, we generate one augmentation of it and encode it by the same multimodal feature encoder to obtain its representation, denoted as $\textbf{x}_{i+}^{S}$.
@@
 The contrastive loss is expressed as follows:
-(3) $\mathcal{L}_{ctr}^{S}=-\sum_{(i,i_{+})\in b}\log\frac{\exp(\mathrm{sim}(\textbf{z}_{i}^{S},\textbf{z}_{i_{+}}^{S})/t)}{\sum_{k=1,k\neq i}^{2|b|}\exp(\mathrm{sim}(\textbf{z}_{i}^{S},\textbf{z}_{k}^{S})/t)}.$
+$\mathcal{L}_{ctr}^{S}=-\sum_{(i,i_{+})\in b}\log\frac{\exp(\mathrm{sim}(\textbf{z}_{i}^{S},\textbf{z}_{i_{+}}^{S})/t)}{\sum_{k=1,k\neq i}^{2|b|}\exp(\mathrm{sim}(\textbf{z}_{i}^{S},\textbf{z}_{k}^{S})/t)}.$ (3)
 Here, $b$ denotes the current batch and $|b|$ denotes the size of the current batch.
@@
 Here, cosine similarity is being used.
-The data augmentation on the target domain training batch and the contrastive loss computation for the target domain $\mathcal{L}_{ctr}^{T}$ is applied in the same way.
-The performance of contrastive learning is largely influenced by the transformation used to generate positive samples (Tian et al., 2020).
-Following the data augmentation experiments in (Bhattacharjee et al., 2023), we include text-level: {synonym replacement, random swap, random crop}, and image-level: {random resize and crop, random horizontal flip, Gaussian blur} into the set of candidate data augmentation techniques (detailed descriptions can be found in the supplementary material).
-We also include the combination of text-level and image-level augmentations as the candidates.
-Our experiments show that all the above strategies result in similar performances (in terms of accuracy on the target domain), but applying Gaussian blur to the image results in the best overall performance (averaged over all domain adaptation settings) on the target domain.
-Thus, in our approach, we apply Gaussian blur to the image while keeping the text unaltered to generate the positive sample of the anchor item.
+The data augmentation on the target domain training batch and the contrastive loss computation for the target domain $\mathcal{L}_{ctr}^{T}$ are applied in the same way.
 Note, the augmentation preserves the label.
-Maximum Mean Discrepancy (MMD): Now that we have learned more separable representations, $\textbf{z}^{S}$ and $\textbf{z}^{T}$, in the projected space, we want to further ensure that these representations capture domain-invariant feature.
-To achieve this, we adopt maximum mean discrepancy (MMD) (Gretton et al., 2012).
-MMD is a kernel-based statistical test used to determine if two samples are drawn from different distributions (Gretton et al., 2012).
+Maximum Mean Discrepancy (MMD): Now that we have learned more separable representations, $\textbf{z}^{S}$ and $\textbf{z}^{T}$, in the projected space, we want to further capture domain-invariant features.
+To achieve this, we adopt maximum mean discrepancy (MMD) Gretton et al.
+(2012).
+MMD is a kernel-based statistical test used to determine if two samples are drawn from different distributions Gretton et al.
+(2012).
 We use it as a loss function to measure the discrepancy between two distributions.
@@
 The squared MMD between $Z^{S}$ and $Z^{T}$ in the reproducing kernel Hilbert space (RKHS) $\mathcal{H}$ is formulated as follows:
-(4) $\mathrm{MMD}^{2}[\mathcal{F},Z^{S},Z^{T}]$ $=\left[\sup_{||f||_{\mathcal{H}}\leq 1}(\mathbf{E}[f(Z^{S})]-\mathbf{E}[f(Z^{T})])\right]^{2}$
+$\mathrm{MMD}^{2}$ $=\left[\sup_{||f||_{\mathcal{H}}\leq 1}(\mathbf{E}[f(Z^{S})]-\mathbf{E}[f(Z^{T})])\right]^{2}$ (4)
 $=||\mu_{S}-\mu_{T}||^{2}_{\mathcal{H}},$
 where $\mu_{S}$ and $\mu_{T}$ are mean embeddings of $f(Z^{S})$ and $f(Z^{T})$, $f\in\mathcal{F}$. $\mathcal{F}$ is the unit ball function class in the Hilbert space $\mathcal{H}$.
-Following prior work (Pan et al., 2010; Long et al., 2015; Bhattacharjee et al., 2023), we compute the MMD between features in the lower dimensional space (i.e.
+Following prior work (Pan et al., 2010; Long et al., 2015; Bhattacharjee et al., 2023), we compute the MMD of features in the lower dimensional space (i.e.
 the projected z space).
-In our implementation, we use the empirical MMD (Gretton et al., 2012), which is expressed as follows:
-(5) $\mathrm{MMD}[\mathcal{F},Z^{S},Z^{T}]=$ $\left[\frac{1}{m^{2}}\sum_{i,j=1}^{m}k(\textbf{z}_{i}^{S},\textbf{z}_{j}^{S})-\frac{2}{mn}\sum_{i,j=1}^{m,n}k(\textbf{z}_{i}^{S},\textbf{z}_{j}^{T})\right.$
+In our implementation, we use the empirical MMD Gretton et al.
+(2012), which is expressed as follows:
+$\mathrm{MMD}=$ $\left[\frac{1}{m^{2}}\sum_{i,j=1}^{m}k(\textbf{z}_{i}^{S},\textbf{z}_{j}^{S})-\frac{2}{mn}\sum_{i,j=1}^{m,n}k(\textbf{z}_{i}^{S},\textbf{z}_{j}^{T})\right.$ (5)
 $\left.+\frac{1}{n^{2}}\sum_{i,j=1}^{n}k(\textbf{z}_{i}^{T},\textbf{z}_{j}^{T})\right]^{\frac{1}{2}}.$
 Here, both $m$ and $n$ are equal to the batch size. $k(\cdot,\cdot)$ denotes the kernel function. $\sigma$ is a free parameter. We use the Radial basis function kernel shown in the following:
-(6) $k(\textbf{z}_{i},\textbf{z}_{j})=\exp\left(-\frac{||\textbf{z}_{i}-\textbf{z}_{j}||^{2}_{2}}{2\sigma^{2}}\right).$
+$k(\textbf{z}_{i},\textbf{z}_{j})=\exp\left(-\frac{||\textbf{z}_{i}-\textbf{z}_{j}||^{2}_{2}}{2\sigma^{2}}\right).$ (6)
 Training: The learned domain-invariant representation, z, is then input to the classifier to get the predicted label $\hat{y}$ for the corresponding out-of-context news:
-(7) $\hat{y}=\mathrm{CLS}(\textbf{z}).$
+$\hat{y}=\mathrm{CLS}(\textbf{z}).$ (7)
 The loss function for the classifier is the cross-entropy loss, where $\hat{y}$ is the predicted label and $y$ is the ground true label:
-(8) $\mathcal{L}_{CE}=-\mathbf{E}_{y\sim Y}\left[y\log(\hat{y})+(1-y)\log(1-\hat{y})\right].$
+$\mathcal{L}_{CE}=-\mathbf{E}_{y\sim Y}\left[y\log(\hat{y})+(1-y)\log(1-\hat{y})\right].$ (8)
 During training, both the projection head and the classifier have learnable parameters.
@@
 This can be expressed as follows:
-(9) $\mathcal{L}=\frac{1}{2}\lambda_{CE}(\mathcal{L}_{CE}^{S}+\mathcal{L}_{CE+}^{S})+\frac{1}{2}\lambda_{ctr}(\mathcal{L}_{ctr}^{S}+\mathcal{L}_{ctr}^{T})+\lambda_{MMD}\mathrm{MMD},$
+$\mathcal{L}=$ $\frac{1}{2}\lambda_{CE}(\mathcal{L}_{CE}^{S}+\mathcal{L}_{CE+}^{S})+\frac{1}{2}\lambda_{ctr}(\mathcal{L}_{ctr}^{S}+\mathcal{L}_{ctr}^{T})$ (9)
+$+\lambda_{MMD}\mathrm{MMD},$
 where $\lambda_{CE}$ is the weight of the cross entropy losses, $\lambda_{ctr}$ is the weight of the contrastive losses and $\lambda_{MMD}$ is the weight of MMD.
-4.3. Test-Time Adaptation
-Domain-invariant features help the model to generalize to the target domain.
+4.3 Test-Time Adaptation
+Domain-invariant features help the model generalize to the target domain.
 In this section, we introduce test-time adaptation which leverages target domain statistics to further adapt the model to the target domain during evaluation.
-Prior work (Li et al., 2016) observes that the statistics of the Batch Normalization (BN) layer contain domain-specific information.
-Concretely, they find that running the test set through the trained model achieves deep adaptation effects on the target domain.
+Prior work Li et al.
+(2016) observes that the statistics of the Batch Normalization (BN) layer contain domain-specific information.
+Concretely, they find that running test set through the trained model achieves deep adaptation effects on the target domain.
 The reason is that this process allows BN statistics to keep track of the target domain statistics, which is conducive to the domain adaptation.
-As such, we introduce BN layers into the classifier.
 At each training step, BN statistics compute and update the running estimates $\{\hat{\mu}_{k+1},\hat{\sigma}^{2}_{k+1}\}$ (Eq.
@@
 (11)) of the layer batch inputs $\{x_{i}\}_{i=1}^{b}$:
-(10) $\hat{\mu}_{k+1}=(1-\rho)\hat{\mu}_{k}+\rho\mu_{k},\quad\hat{\sigma}^{2}_{k+1}=(1-\rho)\hat{\sigma}^{2}_{k}+\rho\sigma^{2}_{k}$
-(11) $\mu_{k}=\frac{1}{b}\sum_{i}x_{i},\quad\sigma^{2}_{k}=\frac{1}{b}\sum_{i}(x_{i}-\mu_{k})^{2}$
+$\hat{\mu}_{k+1}=(1-\rho)\hat{\mu}_{k}+\rho\mu_{k},\quad\hat{\sigma}^{2}_{k+1}=(1-\rho)\hat{\sigma}^{2}_{k}+\rho\sigma^{2}_{k},$ (10)
+$\mu_{k}=\frac{1}{b}\sum_{i}x_{i},\quad\sigma^{2}_{k}=\frac{1}{b}\sum_{i}(x_{i}-\mu_{k})^{2}.$ (11)
 Here, $k$ denotes the $k$-th training step, and $\rho$ is the momentum with a default value of 0.1.
-BN statistics are then fixed (denoted as $\mu$ and $\sigma^{2}$) and used for normalization when the model is set to evaluation mode (pyt, 2024).
-During evaluation, we apply test-time adaptation to calibrate the BN statistics, i.e.
-the mean ($\mu$) and variance ($\sigma^{2}$).
-The BN layer is formulated as follows:
-(12) $x_{out}=\frac{x_{in}-\mu}{\sqrt{\sigma^{2}+\epsilon}}*\gamma+\beta$
-where $x_{in}$ is the BN layer input, $x_{out}$ is the output. $\gamma$ and $\beta$ are BN learnable parameters. $\epsilon$ is added for numerical stability.
-During model training, since only the source domain data is passed through the classifier, BN statistics are computed based on the source domain statistics.
-During model evaluation, the target domain data is passed through the classifier.
-To integrate target domain statistics into the BN statistics, after each training epoch, we pass the unlabeled test set through the classifier.
+$x_{i}$ denotes the input to BN.
+BN statistics are then fixed (denoted as $\mu$ and $\sigma^{2}$) and used for normalization when the model is set to evaluation mode pyt (2024).
+As such, we add BN layers into the classifier.
+To integrate target domain statistics into BN’s memory, we first pass the unlabeled target domain test data through the classifier before setting it to the evaluation mode.
 After that, we set the classifier to evaluation mode and evaluate it on the test set.
 In this way, the running estimates would keep the target domain statistics as well into record, which is helpful for the domain adaptation.
-5. Experimental Design
-5.1. Datasets
+5 Experimental Design
+5.1 Datasets
 We conduct experiments on Twitter-COMMs and NewsCLIPpings, which are the only two out-of-context news datasets that conform to our problem definition (Section 3).
-Twitter-COMMs (Biamby et al., 2022): This dataset is collected from Twitter solely on three topics: Covid-19 (Cv), Climate Change (Cl) and Military Vehicles (M).
-Each original tweet is represented by an image and its associated caption (text).
-Given a caption, the authors create out-of-context news by either randomly choosing an image (random negative) or by choosing the image whose original associated text is semantically the closest (hard negative).
-After running the script provided by the authors to download tweets and removing tweets that has an unavailable image, we obtain 2,143,934 items (Cv: 1,387,043, Cl: 512,490, M: 244,401) in the training set, and 22,082 items (Cv: 6,456, Cl: 8,488, M: 7,138) in the test set.
-The labels have a balanced distribution.
-Hence, we address domain adaptation among different news topics in Twitter-COMMs.
-NewsCLIPpings (Luo et al., 2021): This dataset is derived from the VisualNews dataset (Liu et al., 2020), which is a benchmark news image captioning dataset.
+Twitter-COMMs Biamby et al.
+(2022): This dataset is collected from Twitter solely on three topics: Covid-19 (Cv), Climate Change (Cl) and Military Vehicles (M).
+After downloading the tweets and removing the tweets that has an unavailable image, we obtain 2,143,934 items (Cv: 1,387,043, Cl: 512,490, M: 244,401) in the training set, and 22,082 items (Cv: 6,456, Cl: 8,488, M: 7,138) in the test set.
+The labels are balanced.
+NewsCLIPpings Luo et al.
+(2021): This dataset is derived from the VisualNews dataset (Liu et al., 2020), a benchmark news image captioning dataset.
 The news is collected from four agencies, namely Guardian (G), BBC (B), Washington Post (W) and USA Today (U).
-The NewsCLIPpings dataset is originally partitioned into four splits in which ways the out-of-context news is constructed.
-However, the analysis on the underlying VisualNews dataset (Liu et al., 2020) reveals that there are inter correlations among data from different news agencies (e.g.
-, the image captioning model trained on the USA Today data achieves 3.7 in CIDEr score (Vedantam et al., 2015) when tested on the USA Today test set, yet only achieves 0.6 on the Guardian test set.
-).
-Since NewsCLIPping is derived from VisualNews, we re-partition the dataset according to the news agencies.
 After downloading the dataset, we obtain a total number of 1,182,900 items (B: 205,968, G: 567,012, U: 228,393, W: 181,527) in the training set, and 124,168 items (B: 22,074, G: 59,163, U: 24,243, W: 18,688) in the test set.
-The labels have a balanced distribution.
-Hence, we address domain adaptation among different news agencies in NewsCLIPpings.
-5.2. Implementation Details
-Table 1.
-Evaluation of domain adaptation performance on TwitterCOMMs and NewsCLIPpings.
+The labels are balanced.
+5.2 Implementation Details
+Augmentation: In our approach, we apply Gaussian blur to the image while keeping the text unaltered to generate the augmentation of the anchor item.
+More details and experimental results can be found in Appendix A.1.
+Architecture: We use BLIP-2’s multimodal feature extractor to embed the image and text, and obtain the embedding x of size 768.
+Details of multimodal feature encoder selection can be found in Appendix A.1.
+The projection head is composed of two linear layers with 768 and 500 (experimented with {300, 500, 768}) neurons respectively.
+The classifier is composed of three linear layers with 768, 768 and 2 neurons respectively, and two batch normalization layers following the first two linear layers.
+We use Tanh() as the activation function.
+We apply dropout both before the first and the third linear layer.
+The dropout rate is set to 0.2 (we experiment with {0.2, 0.5} and without dropout).
+Training: We perform training using the Adam optimizer with a batch size of 256 for 20 epochs.
+The early stopping epoch is set to 5.
+The learning rate is set to 2e-4 for Twitter-COMMs and 1e-4 for NewsCLIPpings.
+For the loss function, $\lambda_{MMD}$ is set to 1 and both $\lambda_{CE}$ and $\lambda_{ctr}$ are set to 0.5.
+Source and Target Domain Partition: We treat news toics as individual domains in Twitter-COMMs.
+We use one topic as the target domain and the remaining two topics as the source domain.
+We treat news agencies as individual domains in NewsCLIPpings.
+Based on an observation (Liu et al., 2020) on the correlation of these four news agencies and our pilot experiments, we use two news agencies of the same country as the target domain and the rest two as the source domain respectively.
+5.3 Baseline Models
+Table 1: Evaluation of domain adaptation performance on TwitterCOMMs and NewsCLIPpings.
 Cv, Cl and M stand for covid-19, climate change and military vehicles respectively.
@@
 REAL-FND 73.50 74.82 79.13 79.36 76.64 77.62 69.70 69.96 75.20 75.30 79.95 80.00 78.50 78.52
+CADA 76.17 74.92 79.12 79.11 79.55 79.62 69.67 69.36 75.67 75.86 80.20 80.07 78.28 78.23
 ConDA-TTA 79.26 79.34 80.80 80.80 80.06 80.08 71.52 71.80 75.70 75.86 80.77 80.84 77.17 77.31
-Multimodal Feature Encoder Selection: To find the optimal MLLM feature extractor, we experiment with (i) BLIP-2 (Li et al., 2023b) multimodal and (ii) unimodal feature extractor, (iii) CLIP (Radford et al., 2021) (only has unimodal feature extractors) and (iv) ALBEF (Li et al., 2021b) multimodal feature extractor.
-For (ii) and (iii), we experiment with vector concatenation and element-wise multiplication to obtain the multimodal feature vector.
-We then train a one-layer linear classifier on the four features respectively and evaluate the performances.
-Experimental results show that BLIP-2 multimodal features gives the best performance (in terms of accuracy) on both the training and test sets.
-Note, we also evaluate the zero-shot performance of the BLIP-2 Image Text Matching head which only outputs a similarity score rather than the feature vector.
-We find it underperforms (around 0.5 in accuracy on both training and test set) using BLIP-2 multimodal feature directly.
-Architecture: We use BLIP-2’s blip2_feature_extractor with the multimodal mode to embed the image and text, and obtain the multimodal embedding x of size 768.
-The projection head is composed of two linear layers with 768 and 500 (experimented with {300, 500, 768}) neurons respectively.
-The classifier is composed of three linear layers with 768, 768 and 2 neurons respectively, and two batch normalization layers following the first two linear layers.
-We use Tanh() as the activation function.
-We apply dropout both before the first linear layer and the third linear layer.
-The dropout rate is set to 0.2 (we experimented with {0.2, 0.5} and without dropout).
-Training: We perform training using the Adam optimizer with a batch size of 256 for 20 epochs.
-The early stopping epoch is set to 5.
-The learning rate is set to 2e-4 for the Twitter-COMMs dataset and 1e-4 for the NewsCLIPpings dataset.
-For the loss function, $\lambda_{MMD}$ is set to 1 and both $\lambda_{CE}$ and $\lambda_{ctr}$ are set to 0.5.
-[…]
-5.3. Baseline Models
 Considering that no prior work has looked into domain adaptive out-of-context news detection, we select SOTA baseline models from a similar task: domain adaptive fake news detection.
-For fair comparison, we use BLIP-2’s blip2_feature_extractor with the multimodal mode to encode the image and text into a multimodal feature vector.
+For fair comparison, we use the same multimodal feature encoder for all baselines.
 Source only serves as the naive baseline for the chosen backbone feature encoder. It uses one linear layer and the softmax function as the classifier. During training, it only uses the source domain.
@@
 Therefore, it can be viewed as the upper bound performance.
-EANN (Wang et al., 2018) is a multimodal fake news detection model that uses adversarial learning to derive event-invariant features which benefit fake news detection on newly arrived events.
-Since the original EANN implementation encodes the image and text modality separately, we modify it to accommodate multimodal inputs.
-MDA-WS (Li et al., 2021a) is an early fake news detection framework based on multi-source domain adaptation and weak supervision.
-It uses adversarial learning to learn the domain-invariant feature, and incorporates prior knowledge to assign pseudo labels to target domain news for weak supervision.
-CANMD (Yue et al., 2022) is a domain adaptive fake news detection model.
-It firstly pre-trains the model on the source training data and then uses the pseudo labels generated from it on the target training data to adapt the model.
-During adaptation, it aims to reduce intra-class discrepancy and enlarge inter-class discrepancy.
-REAL-FND (Mosallanezhad et al., 2022) is a reinforcement-learning (RL) based domain adaptive fake news detection model.
-It first obtains a reasonable news representation and then trains the fake news and domain classifiers.
-Using the representation as the state and the classifiers’ losses as the reward, it then trains an RL agent that learns a domain-invariant news representation.
-6. Result and Analysis
+EANN Wang et al. (2018) is a multimodal fake news detection model that uses adversarial learning to derive event-invariant features which benefit fake news detection on newly arrived events.
+MDA-WS Li et al. (2021b) uses adversarial learning to learn the domain-invariant feature, and incorporates prior knowledge to assign pseudo labels to target domain news for weak supervision.
+CANMD Yue et al. (2022) firstly pre-trains the model on the source training data and then generates pseudo labels for the target training data to adapt the model.
+REAL-FND Mosallanezhad et al. (2022) uses news representation as the state and the classifiers’ losses as the reward, and trains an RL agent that learns a domain-invariant news representation.
+CADA Li et al. (2023a) proposes class-based adversarial domain adaptive framework to achieve fine-grained alignment.
+6 Result and Analysis
 Based on the above experimental setup, we evaluate our proposed ConDA-TTA and investigate the contribution of each model component afterwards.
-We also include TSNE visualization and parameter sensitivity analysis.
-6.1. Experimental Result
-Table 2.
-Evaluation results on ablating $\mathcal{L}_{ctr}$, MMD and TTA of ConDA-TTA.
-Cv, Cl and M stand for covid-19, climate change and military vehicles respectively.
-B, G, U and W stand for BBC, Guardian, USA Today and Washington Post respectively.
-In X $\rightarrow$ Y, X denotes the source domain, Y denotes the target domain.
-The best performance is in bold text.
-The second best performance is underlined.
+We also include TSNE visualization and parameter sensitivity analysis (Appendix A.3).
+6.1 Experimental Result
+Table 1 shows the evaluation performance of ConDA-TTA and the baseline models. We use F1 (%) and accuracy (Acc%) as the evaluation metrics as in Luo et al. (2021) and Biamby et al. (2022).
+Naive Baseline and Upper Bound: We first present our observations on the naive baseline (source only) and the supervised training (source and target) results.
+In our experiments, we find that using BLIP-2’s multimodal representation with a plain linear classifier gives competitive performances.
+However, we observe around 2-10% performance gap between the naive baseline and the supervised training.
+This suggests the need for developing domain adaptation techniques to mitigate the gap.
+The domain adaptation gaps of different settings vary.
+For news topics in Twitter-COMMs, adapting to Military Vehicles (M) has the largest domain adaptation gap (10.6% in F1), whilst adapting to Climate Change (Cl) and Covid-19 (Cv) have relatively smaller domain adaptation gaps.
+We conjecture that the differences are potentially caused by the different news styles.
+We manually check a batch of 10 randomly sampled news from each topic.
+We find that Cl and Cv tend to include images that have more abundant information, containing such as embedded texts, numbers, and even news screenshots.
+However, in M, the news image usually contains one clear subject: the vehicle itself.
+At the same time, it requires more specialized knowledge to determine whether the image pairs the text within this topic.
+For news agencies in NewsCLIPpings, adapting to BBC (B) and USA Today (U) have a larger domain adaptation gap (6.81% and 4.43% in accuracy), while adapting to the Guardian (G) and the Washington Post (W) have a relatively smaller gap.
+Although the gap is not consistently large, it could cause a big difference on news platforms where a large amount of news is posted everyday.
+Thus, we believe that it is crucial to leverage MLLM feature encoders and further mitigating its domain adaptation gap with a small training cost.
+Baseline Comparison: We next compare our model’s performance with the baseline models.
+On both datasets, our proposed ConDA-TTA outperforms the baseline models across most metrics.
+Notably, on Twitter-COMMs, ConDA-TTA outperforms the best baseline model by 2.93% in F1 when the source domains are Covid-19 (Cv) and Climate Change (Cl), and the target domain is Military Vehicles (M).
+On NewsCLIPpings, when the target domain is BBC (B) and USA Today (U), our model outperforms the baselines.
+When the target domain is Guardian (G), our model achieves the second best performance.
+However, we observe that when the target domain is the Washington Post (W), all models see slightly negative transfer.
+This is in line with the observations on the underlying dataset Liu et al.
+(2020), where the authors find that training their image captioning model on either the B or G results in a comparatively low CIDEr score (Vedantam et al., 2015) on the W.
+As such, we conjecture that the information contained in B and G is not enough for enhancing the model’s domain adaptation performance on W.
+Table 2: Evaluation results on ablating $\mathcal{L}_{ctr}$, MMD and TTA of ConDA-TTA.
 Model Twitter-COMMs NewsCLIPpings
@@
 ConDA-TTA 79.26 79.34 80.80 80.80 80.06 80.08 71.52 71.80 75.70 75.86 80.77 80.84 77.17 77.31
-Table 1 shows the evaluation performance of ConDA-TTA and the baseline models. We use F1 and accuracy (Acc) as the evaluation metrics.
-Naive Baseline and Upper Bound: We first present our observations on the naive baseline (source only in Table 1) and the supervised training (source and target in Table 1) results.
-In our experiments, we find that using BLIP-2’s multimodal representation with a plain linear classifier already achieves strong performances (around 70 - 80% zero-shot accuracy under most domain adaptation scenarios) on this task.
-Nevertheless, we observe around 2-10% performance gap between the naive baseline and the supervised training result.
-Considering that BLIP-2 has 2.7b parameters, it is not trivial to apply fine-tuning to improve its domain adaptation performance.
-This suggests the need for developing domain adaptation techniques on top of it to mitigate the gap.
-For the news topics in Twitter-COMMs, adapting to Military Vehicles (see Section 5.1) has the largest domain adaptation gap (10.6% in F1 and 8.32% in accuracy), whilst adapting to Climate Change and Covid-19 has a relatively smaller domain adaptation gap (2.99% in F1 and 2.65% in accuracy, 4.58% in F1 and 3.71% in accuracy).
-We conjecture that the differences are potentially caused by the different news styles.
-We manually check a batch of 10 randomly sampled news from each topic.
-We find that Climate Change and Covid-19 tend to include images with more abundant information, including things such as embedded texts, numbers, and even news screenshots.
-However, for the Military Vehicles domain, the news’ image usually contains one clear subject: the military vehicle itself.
-At the same time, it requires more specialized knowledge to determine whether the image pairs the text within this topic.
-For the news agencies in NewsCLIPpings, adapting to BBC and USA Today has the largest domain adaptation gap (6.61% in F1 and 6.81% in accuracy, 2.93% in F1 and 4.43% in accuracy), while adapting to the Guardian and the Washington Post has a relatively smaller gap.
-Although the gap is not consistently significant, it can affect heavily if deployed in a news platforms where a large amount of news is posted everyday.
-Thus, we believe that leveraging these powerful MLLM feature encoders and further mitigating its domain adaptation gap with a small training cost are crucial to achieve promising performance on unlabeled, newly emergent domains.
-Baseline Comparison: We next compare our model’s performance with the baseline models.
-On both datasets, our proposed ConDA-TTA outperforms the baseline models across most metrics.
-Notably, on the Twitter-COMMs dataset, ConDA-TTA outperforms the best baseline model by 2.93% in F1 and by 2.08% in accuracy when the source domains are Covid-19 (Cv) and Climate Change (Cl), and the target domain is Military Vehicles (M).
-Specifically, we observe negative transfer for CANMD under this setting.
-Moreover, ConDA-TTA achieves 80.80% on both F1 and accuracy when the source domains are Covid-19 and Military Vehicles, and the target domain is Climate Change, surpassing the baseline performances by 0.93% and 0.27% respectively.
-Additionally, it achieves 80.06% and 80.08% on F1 and accuracy when the source domains are Climate Change and Military Vehicles, and the target domain is Covid-19.
-On the NewsCLIPpings dataset, when the source domains are USA Today (U) and Washington Post (W) and the target domain is BBC (B), our model outperforms the best baseline model by 1.84% in accuracy and 1.82% F1 accuracy.
-When the source domains are BBC (B) and Guardian (G) and the target domain is USA Today (U), our model outperforms the best baseline model by 0.82% in F1 and 0.84% in accuracy.
-When the source domains are USA Today and Washington Post, and the target domain is Guardian, our model achieves the second best performance in both metrics.
-However, we also observe that when the source domains are BBC and Guardian, and target domain is Washington Post, all models see slightly negative transfer in terms of accuracy.
-This is in-line with the observations (Section  5.1) on the underlying dataset (Liu et al., 2020), where the authors find that training their image captioning model on either the BBC or Guardian results in a comparatively low CIDEr score on the Washington Post.
-As such, we conjecture that the information contained in the BBC and Guardian is not enough for enhancing the model’s domain adaptation performance on the Washington Post.
 […]
-6.2. Ablation Study
-We next present an ablation study on the different components in ConDA-TTA. Specifically, we ablate $\mathcal{L}_{ctr}$, MMD and TTA from ConDA-TTA. The evaluation results are summarized in Table 2.
-On the Twitter-COMMs dataset, we observe that ablating MMD causes the biggest performance drop in all three domain adaptation settings, respectively: 3.41% (F1) and 3.42% (Acc), 0.33% (F1) and 0.32% (Acc), 0.92% (F1) and 0.91% (Acc).
-This suggests that MMD is the most important component in ConDA-TTA.
-Additionally, ablating the contrastive loss results in the second biggest performance decrease when the target domain is Military Vehicles and Climate Change, respectively: 1.19% (F1) and 1.22% (Acc), 0.21% (F1) and 0.21% (Acc).
-However, we see a marginal performance increase (0.04% in F1 and 0.03% in Acc) when the target domain is Covid-19.
-Lastly, ablating TTA causes a 0.68% (F1) and 0.75% (Acc) performance drop when the target domain is Military Vehicles, and causes a 0.07% performance drop in both F1 and Acc when the target domain is Climate Change.
-When the target domain is Covid-19, we again see a slight increase (0.46% in F1 and 0.50% in Acc) in performance.
-Although ConDA-TTA (as a whole) improves the domain adaptation performance on Covid-19 (Section 6.1), not every component of it contributes under this setting.
-On the NewsCLIPpings dataset, we observe similar degradation when ablating these three components when the target domain is the BBC, Guardian and USA Today.
-Specifically, when ablating TTA, the performance decreases (1.22% in F1 ad 1.53% in Acc) most significantly on BBC, while less significantly on Guardian (0.02% in F1 and 0.18% in Acc) and USA Today (0.02% in F1 and 0.07% in Acc).
-When ablating MMD, F1 drops by 0.04%, 0.06% and 0.84%, and the Acc drops by 0.04%, 0.06% and 0.86%.
-When ablating the contrastive loss, the F1 drops by 0.62%, 0.63% and 0.5%, and the Acc drops by 0.53%, 0.58% and 0.51%.
-However, when adapting to Washington Post, ablating the contrastive loss, MMD and TTA all bring slight performance increase (0.04% - 0.72%).
-We conjecture that this observation echos previous analysis (Section 6.1) that the information contained in BBC and Guardian is not enough for assisting the domain adaptation on Washington Post, making each component less effective.
-In summary, the above results confirm the efficacy of the contrastive loss, MMD and TTA in most domain adaptation scenarios on both datasets.
-Concretely, MMD tends to contribute the most to the domain adaptation on Twitter-COMMs; TTA tends to contribute the most on NewsCLIPpings.
-However, we note that, in some cases, the techniques are less effective on the NewsCLIPpings dataset.
-We believe this is due to that under some settings, the source domain contains less abundant information for adaptation to the target domain (Section 6.1).
-6.3. Visualization
-To show that ConDA-TTA effectively learns the domain-invariant feature, we adopt TSNE (Van der Maaten and Hinton, 2008) to visualize the multimodal embeddings x (also the input embedding to ConDA-TTA) and the learned domain-invariant feature z projected to the 2-d space (Figure 3).
-Figure 3. TSNE visualization of the multimodal feature x and the learned domain-invariant feature z under Cv, Cl $\rightarrow$ M.
-Here we illustrate the TSNE visualization of setting Cv, Cl $\rightarrow$ M.
-We can tell from the figure that feature z from different topics blend better than x, where we see clearer boundaries between different topics.
-This suggests that our model has removed domain-specific features from x to some extent.
-This further shows that our approach is able to effectively learn domain-invariant feature, and transfers out-of-context news detection knowledge to the target domain (Table 1).
-6.4. Sensitivity Analysis
-The contrastive loss and the MMD are two crucial components in our approach.
-This section briefly studies how their weights, $\lambda_{ctr}$ and $\lambda_{MMD}$, in the loss function affect the performance of ConDA-TTA.
-Figure 4. ConDA-TTA’s performance (in Acc) with different $\lambda_{MMD}$ and $\lambda_{ctr}$ values. The legend denotes the target domain.
-Figure 4 illustrates ConDA-TTA’s performance with $\lambda_{MMD}=\{0.5,1,2,5\}$ and $\lambda_{ctr}=\{0.1,0.5,1,2\}$.
+6.2 Ablation Study
+We next present an ablation study on different components in ConDA-TTA. Specifically, we ablate $\mathcal{L}_{ctr}$, MMD and TTA from ConDA-TTA. The evaluation results are summarized in Table 2.
+On Twitter-COMMs, we observe that ablating MMD causes the biggest performance drop in all three domain adaptation settings.
+Additionally, ablating $\mathcal{L}_{ctr}$ and TTA both result in performance decrease when the target domain is Military Vehicles (M) and Climate Change (Cl).
+However, we see a marginal performance increase when the target domain is Covid-19 (Cv).
+On NewsCLIPpings, we observe similar degradation when ablating these three components for target domain being BBC (B), Guardian (G) and USA Today (U).
+Specifically, when ablating TTA, the performance decreases most significantly on B, and less significantly on G and U.
+However, when adapting to Washington Post (W), ablating $\mathcal{L}_{ctr}$, MMD and TTA all bring slight performance increase.
+We assume that this observation echos the previous analysis (Section 6.1) that the information contained in BBC and Guardian is not enough for helping the domain adaptation to Washington Post, therefore making each component less effective.
+In summary, the above results confirm the efficacy of $\mathcal{L}_{ctr}$, MMD and TTA under most domain adaptation scenarios on both datasets.
+Concretely, MMD tends to contribute the most on Twitter-COMMs; TTA tends to contribute the most on NewsCLIPpings.
+It is worth noting that news in Twitter-COMMs are more colloquial and informal, leading writing styles vary much across different topics.
+We conclude that MMD is better at capturing domain-invariant features in this scenario.
+6.3 Visualization
+To show that ConDA-TTA effectively learns the domain-invariant feature, we adopt TSNE Van der Maaten and Hinton (2008) to visualize the multimodal feature x and the learned domain-invariant feature z projected to the 2-d space (Figure 3).
+Figure 3: TSNE visualization of the multimodal feature x and the learned domain-invariant feature z under Cv, Cl $\rightarrow$ M.
+We can tell from the figure that feature z of different topics blend better than x.
+Additionally, in Table 3, we compute the variance of x and z after projecting and normalizing them onto 1-d space.
+The decrease of variance and the TSNE visualization suggest that our model has removed domain-specific features from x to some extent.
+This further shows that our approach effectively learns domain-invariant features.
+The parameter sensitivity analysis can be found in Appendix A.3.
+Table 3: Variance of x and z on different domains.
+Feature Var(M) Var(Cl) Var(Cv)
+x 0.0623 0.0738 0.0780
+z 0.0326 0.0652 0.0724
+7 Conclusion
+This paper has proposed ConDA-TTA, a domain adaptive out-of-context news detection model that can adapt to both unlabeled news topics and news agencies.
+ConDA-TTA first encodes the news using the BLIP-2 multimodal feature encoder and then adopts contrastive learning and MMD to learn domain-invariant features.
+During test-time, it further incorporates target domain statistics into the classifier.
+Experimental results show the effectiveness and superiority of our approach.
+Limitations
+Despite the effectiveness of our proposed ConDA-TTA model, there are several limitations.
+First, in this work, we focus on two domain types: news topics and news agencies.
+We do not consider other potential types of domain, such as news from different regions, languages, or cultural backgrounds.
+Second, we observe that under some domain adaptation settings, all experimented models exhibit negative transfer.
+In this paper, we present analysis based on the findings in the underlying dataset paper.
+We will leave a deeper analysis into our future work.
+Appendix A Appendix
+A.1 Implementation Details
+Augmentation: The performance of contrastive learning is largely influenced by the transformation used to generate positive samples (Tian et al., 2020).
+Following the data augmentation selection in  Bhattacharjee et al.
+(2023), we experiment with a set of data augmentation techniques and select the one with the best performance on the test set as the final strategy.
+In view of the large size of the training set (Section 5.1), we randomly sample 1.5% data items from the training set as a toy training set, roughly the same size as the test set.
+We do this in order to expedite the selection of positive sample generation strategy.
+We experiment with a set of augmentations: text-level: {synonym replacement, random swap, random crop}, image-level: {random resize and crop, random horizontal flip, Gaussian blur}, and the combination of text-level and image-level augmentations.
+Below is the detailed descriptions of them:
+Text-level:
+  • Synonym Replacement: Following Bhattacharjee et al.
+(2023), only words identified as nouns, adjectives, adverbs and verbs will be considered for synonym replacement.
+We randomly choose $n$ words of these parts-of-speech to be replaced, where $n=10\%\times$ the number of words in the sentence.
+We use nltk.corpus.wordnet to obtain the synonyms and randomly choose one for each original word.
+  • Random Swap: We randomly swap $n$ pair of words in the sentence, where $n=10\%\times$ the number of words in the sentence. This is achieved by using the nlpaug package Ma (2019).
+  • Random Crop: We randomly remove $n$ consecutive words from the sentence, where $n=10\%\times$ the number of words in the sentence. This is achieved by using the nlpaug package Ma (2019).
+Image-level:
+  • Random Resize and Crop: We crop a random region of the image resize it to (224, 224).
+  • Random Horizontal Flip: We flip the image horizontally with a probability of 0.5.
+  • Gaussian Blur: We blur the image with the random Gaussian kernel.
+The kernel size range is from 5 to 9.
+The standard deviation range is from 0.1 to 5.0.
+All of the image-level augmentations are achieved by using torchvision.transforms.
+Our experiments (Table 4) show that Gaussian blur results in the best overall performance.
+Thus, in our approach, we apply Gaussian blur to the image while keeping the text unaltered to generate the positive sample of the anchor item.
+Table 4: Evaluation on different augmentation strategies of ConDA-TTA on Twitter-COMMs. Performances are reported in accuracy. The best performance is in bold text.
+Augmentation Target Domain
+M Cl Cv Avg
+Synonym Replacement 79.7 80.2 80.3 80.07
+Random Swap 79.9 80.2 80.3 80.00
+Random Crop 79.3 80.5 80.0 79.93
+Random Resize and Crop 80.1 80.2 79.9 80.07
+Random Horizontal Flip 79.4 80.7 80.0 80.03
+Gaussian Blur 80.0 80.4 80.1 80.16
+Synonym Replacement 79.5 80.4 80.5 80.13
++ Gaussian Blur
+Multimodal Feature Encoder Selection: We experiment with (i) BLIP-2 (Li et al., 2023c) multimodal and (ii) unimodal feature extractor, (iii) CLIP unimodal feature extractor and (iv) ALBEF (Li et al., 2021a) multimodal feature extractor.
+For (ii) and (iii), we experiment with vector concatenation and element-wise multiplication to obtain the multimodal feature.
+We then train a one-layer linear classifier with the four features as inputs respectively and evaluate the performances (Table 5).
+Experimental results show that BLIP-2 multimodal features gives the best performance (in terms of accuracy) on both the training and test sets, as well as having the significant domain adaptation gap.
+We also evaluate the zero-shot performance of the BLIP-2 Image Text Matching head which only outputs a similarity score rather than the feature vector.
+We find it underperforms (achieving around 0.5 in accuracy on both training and test set) using BLIP-2 multimodal feature directly.
+Table 5: Evaluation on different multimodal feature encoders of ConDA-TTA on Twitter-COMMs. Performances are reported in accuracy.
+Multimodal Feature Encoder Target Domain
+M Cl Cv
+BLIP-2multi (supervised) 80.7 81.5 81.6
+BLIP-2multi (source only) 72.4 78.8 77.9
+$\Delta$ domain adaptation gap 8.3 2.7 3.7
+ALBEF (supervised) 69.3 66.8 67.3
+ALBEF (source only) 63.7 64.3 66.0
+$\Delta$ domain adaptation gap 5.6 2.5 1.3
+CLIP (supervised) 76.6 79.3 78.6
+CLIP (source only) 72.1 77.3 78.1
+$\Delta$ domain adaptation gap 4.5 2.0 0.5
+BLIP-2uni (supervised) 72.5 75.0 74.4
+BLIP-2uni (source only) 58.8 71.3 72.1
+$\Delta$ domain adaptation gap 13.7 3.7 2.3
+A.2 Computing Infrastructure
+Our experiments are conducted on one NVIDIA A40 GPU. Each epoch takes around 20 mins to finish.
+A.3 Sensitivity Analysis
+The contrastive loss, MMD and TTA are three crucial components in our approach.
+This section briefly studies how the weights, $\lambda_{ctr}$ and $\lambda_{MMD}$, in the loss function and the batch size $b$ in TTA affect the performance of ConDA-TTA.
+Figure 4: ConDA-TTA’s performances (in Acc) with different $\lambda_{MMD}$ and $\lambda_{ctr}$ values. The legend shows the target domain.
+Figure 5: ConDA-TTA’s performances (in Acc) with different batch sizes. The legend shows the target domain.
+Figure 4 illustrates ConDA-TTA’s performances with $\lambda_{MMD}=\{0.5,1,2,5\}$ and $\lambda_{ctr}=\{0.1,0.5,1,2\}$.
 When setting $\lambda_{MMD}$ to 1, our model achieves the best performance in 4 (M, Cl, B, U) out of 7 domain adaptation settings, and achieves close to best performance in the remaining 3 settings.
 Overall, $\lambda_{MMD}$ being too high causes the accuracy to drop a little.
-This might be because the model has put comparatively less emphasis on learning the detection of out-of-context news (i.e.
-the cross-entropy loss in Eq.
-(9) receives less attention).
+This might be because the model has put relatively less emphasis on the cross-entropy losses in Eq.
+(9).
 For $\lambda_{ctr}$, we find that setting it to 0.5 achieves the best performance in 5 settings.
-When the value goes up, the model performance tends to decrease gradually.
-We conjecture this is because the similarity (Section 4.2) among original news becomes smaller, making it slightly difficult to find the decision boundary.
-7. Conclusion
-This paper has proposed ConDA-TTA, a domain adaptive out-of-context news detection model that can adapt to both unlabeled news topics and news agencies.
-ConDA-TTA first encodes the news using the BLIP-2 multimodal feature encoder and then adopts contrastive learning and MMD to learn the domain-invariant features.
-During test-time, we further incorporate target domain statistics into the classifier of ConDA-TTA, in order to better adapt the model to the target domain.
-Experimental results show the efficacy of our approach.
-These advantages enable ConDA-TTA to transfer the learned out-of-context news detection knowledge to emerging events or news platforms, lessening the burden of online news content moderators.
+we find that when its value goes up, the model performance tends to decrease gradually.
+We conjecture this is because the distances (Section 4.2) between news with the same label have become much larger, making it more difficult to classify.
+Figure 5 illustrates ConDA-TTA’s performances with test set’s batch size $b=\{64,128,256,512\}$.
+Overall, if computing resources permit, larger batch size tend to result in better model performances on most domain adaptation settings.
+However, the differences are not much.
+We argue that larger batch size helps the TTA to learn a less biased target domain statistics, because it includes more data points in one batch.
+However, we also conclude that it’s influence is limited.